### v1

# NestedFP: High-Performance, Memory-Efficient Dual-Precision Floating Point Support for LLMs

Abstract
Large Language Models (LLMs) are playing a crucial role in latency-critical, high-throughput services like virtual assistants and code generation. While techniques such as continuous batching and paged attention address service-level objectives (SLOs), and quantization methods accelerate inference, the dynamic and efficient adaptation of precision at runtime remains a significant, largely underexplored challenge. The emergence of hardware support for FP8 arithmetic, offering up to 2$\times$ the throughput of FP16, presents an attractive opportunity for interactive LLM serving. However, current approaches like co-deploying FP8 and FP16 models suffer from increased storage overhead and fail to unlock FP8’s full potential. To address these limitations, we introduce NestedFP, a novel precision-adaptive serving technique enabling seamless FP8 and FP16 inference from a single 16-bit model representation, thereby incurring no additional memory cost. NestedFP decomposes each FP16 weight into two 8-bit components, facilitating efficient FP8 execution while preserving full FP16 accuracy. We demonstrate the practical viability of our approach by implementing a custom CUTLASS-based GEMM kernel that reconstructs FP16 operands on-the-fly, integrated within the vLLM serving framework. Our evaluation shows that NestedFP delivers up to 1.55$\times$ throughput improvement in FP8 mode with negligible accuracy degradation compared to FP16 precision, while introducing only 3.9% performance overhead on average in FP16 mode across various models. NestedFP thus provides a flexible foundation for dynamic, SLO-aware precision selection, paving the way for more scalable and efficient LLM serving under bursty and heterogeneous workloads.
1 Introduction
Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information. Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand. For example, OpenAI reportedly generates over a billion words per day—equivalent to more than one million words per second marlin —with production query rates frequently surpassing 200 requests per second scale_llm . This widespread deployment in latency-sensitive, user-facing services has intensified the need for inference infrastructures that are not only scalable and efficient, but also capable of satisfying stringent latency and throughput constraints.
The latency of LLM inference is commonly evaluated using two standard metrics: Time To First Token (TTFT), which measures the prefill latency, and Time Per Output Token (TPOT), which captures the average latency for generating each output token. For interactive applications, industry-standard service-level objectives (SLOs) typically require TTFT to be under 200 ms and TPOT to remain below 33.3 ms in order to ensure a responsive user experience baseten_slo . To meet these stringent latency targets under growing demand, several optimization techniques have been proposed. ORCA orca introduces efficient batching mechanisms that consolidate linear layer execution across both the prefill and decode phases. PagedAttention paged_attention addresses memory inefficiencies arising from unpredictable output sequence lengths by reducing over-allocation, thereby increasing the number of concurrently served requests. More recently, Sarathi-Serve sarathi_serve proposes chunked execution of the prefill phase and constructs hybrid batches combining prefill and decode requests to further improve scheduling efficiency.
Recent studies have shown that the FP8 format can achieve accuracy comparable to FP16 across a variety of large language model tasks fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 . The introduction of native FP8 support on modern accelerators—such as NVIDIA Hopper GPUs and Intel Gaudi HPUs—has made FP8 increasingly practical for large-scale inference. On NVIDIA Hopper, FP8-compatible tensor cores can offer up to 2× higher peak throughput compared to their FP16 counterparts, making FP8 particularly advantageous for latency-sensitive or bursty workloads. In this work, we propose leveraging both FP8 and FP16 representations to better satisfy service-level objectives (SLOs). However, co-deploying models in both formats is non-trivial, as it is difficult to ensure robust accuracy and performance without introducing additional storage overhead beyond the originally required 16-bit model.
We present NestedFP, a technique that enables dynamic selection between FP8 and FP16 precision from a single, unified model representation. By analyzing the value distribution of FP16 weights, our approach restructures each weight tensor into two 8-bit components, enabling accurate FP16 reconstruction while supporting efficient FP8 execution—all with no additional memory storage. To realize this format, we develop a custom General Matrix Multiplication (GEMM) kernel built on the CUTLASS library cutlass , which reconstructs FP16 values from the two 8-bit components on-the-fly during GEMM kernel execution. The kernel incurs an average performance overhead of 6.47% across varying GEMM shapes, relative to the vanilla CUTLASS kernels, respectively. Moreover, our 8-bit tensor for FP8 inference achieves FP8 inference accuracy comparable to that of standard FP8 quantized models.
We integrate our custom kernel into the vLLM serving framework vllm and show that it incurs only 3.90% end-to-end performance overhead on average across models in FP16 mode, while delivering up to 1.55$\times$ performance improvements when executing in FP8 mode. NestedFP provides a flexible and memory-efficient foundation for adaptive precision serving, making it well-suited for fluctuating request rates and stringent latency constraints.
2 Background
2.1 Floating Point Representation
[…]
$X_{FP}=(-1)^{S}\cdot\left(1+\sum_{j=1}^{y}M_{j}\cdot 2^{-j}\right)\cdot 2^{E-b},\quad E\in\{0,1,\ldots,2^{x}-1\}$ (1)
Commonly used 16-bit formats include FP16 (E5M10) and BF16 (E8M7), while prevalent 8-bit formats include E4M3 and E5M2. These formats are natively supported by many modern hardware platforms and have become standard choices in low-precision machine learning computations investigation_fp8_across_accelerators .
2.2 FP8 Quantization for LLMs
Rise of FP8 Quantization for LLMs. While FP16 has become the de facto standard for LLM serving, 8-bit floating-point formats (FP8) are gaining traction for scenarios demanding even greater efficiency. This trend is driven by two main factors: (1) the representational advantages of floating-point formats, and (2) the increasing hardware support for FP8 arithmetic.
First, floating-point quantization is generally better suited for LLMs than integer quantization due to its wider dynamic range. Unlike in computer vision tasks—where integer quantization is prevalent—LLMs often exhibit activation outliers amplified by operations such as LayerNorm efficient_ptq_fp8 . These outliers can significantly degrade model performance if not properly managed investigation_fp8_across_accelerators . Although recent work has attempted to mitigate this issue in integer quantization through static outlier-aware techniques efficient_ptq_fp8 ; smoothquant , these methods still underperform compared to floating-point quantization, which inherently better accommodates long-tailed distributions and high dynamic ranges llm_fp4 ; fp8_formats_for_deep_learning ; int_or_fp .
Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs investigation_fp8_across_accelerators . Leveraging this support, FP8 GEMM operations can achieve up to 2×2$\times$2 × speedup compared to their FP16 counterparts. This growing availability is accelerating the adoption of floating-point quantization—particularly FP8—further highlighting its relevance for efficient LLM deployment.
E4M3 Format. Among the two widely adopted FP8 formats—E4M3 and E5M2—existing studies have consistently shown that E4M3 yields higher inference accuracy for Transformer-based language models fp8_formats_for_deep_learning ; efficient_ptq_fp8 ; hybrid_fp8_training_and_inference . However, due to its reduced dynamic range compared to FP16, the effectiveness of E4M3 critically depends on the application of appropriate scaling strategies the_power_of_exponent ; fp8_formats_for_deep_learning ; hybrid_fp8_training_and_inference . Scaling is applied to align the dynamic range of tensor values with the numerical bounds of the E4M3 format. Scaling factors may be determined statically—based on heuristics such as the absolute maximum, percentile clipping, or mean squared error (MSE) minimization fp8_formats_for_deep_learning —or computed dynamically during runtime. Weight tensors are typically scaled statically on a per-channel basis llm_fp4 ; efficient_ptq_fp8 ; int_or_fp , most commonly using the absolute maximum value. Activation tensors, by contrast, may be scaled offline on a per-tensor basis llm_fp4 ; efficient_ptq_fp8 ; int_or_fp or dynamically on a per-token basis during inference. While the latter approach can improve accuracy, it incurs additional runtime cost due to the overhead of computing scaling factors on-the-fly investigation_fp8_across_accelerators ; training_and_inference_llm_fp8 .
FP8 Results. While FP8 quantization offers significant efficiency benefits and shows promising results, it still lags behind FP16 in terms of accuracy. Table 1 presents a comparison between FP8 and FP16 across various benchmarks that require sophisticated reasoning capabilities. The FP8 results are obtained using the vLLM framework vllm , with E4M3 quantization applied using per-channel weight scaling and per-token activation scaling based on absolute maximum value. Quantization is applied exclusively to linear layers, with embedding layers left in higher precision. FP8 generally exhibits a measurable drop in accuracy, highlighting its current limitations for tasks that demand high output fidelity.
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ Minerva Math 18.1 17.6 --0.54 17.1 16.9 --0.20 42.7 42.9 $+$0.16 35.4 34.6 --0.84 MMLU Pro 33.3 32.8 --0.49 35.4 34.9 --0.44 53.1 52.7 --0.37 54.3 53.6 --0.71 BBH 39.5 38.6 --0.83 41.7 41.2 --0.49 27.5 27.1 --0.38 52.5 51.8 --0.65
Table 1: Downstream task accuracy for FP8 quantized models relative to FP16 models.
3 Motivation
3.1 Challenge: Dynamic Resource Demand in LLM Serving
LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests andes ; hygen . These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in TTFT and TPOT. Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet service-level objectives (SLOs).
To quantify the extent of this variability, we analyzed production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace . These traces contain timestamps of incoming requests, from which we constructed the per-second request rate on May 10, 2024, as shown in Figure 1. Our analysis reveals that the request rate ranged from 0 to 100 requests/s across the entire day, and exhibited intra-day variations of up to 5.8-fold within the most variable 1-hour window (14:00–15:00 UTC; min: 17, max: 98) and 3.2-fold within the most variable 1-minute window (18:12–18:13 UTC; min: 31, max: 98).
To cope with such highly variable request patterns, auto scaling is a commonly used to automatically adjusts computational resources based on real-time demand auto_scaling . However, auto scaling is feasible mainly in cloud environments with sufficient available resources, and is typically limited to time scales of few minutes to hours auto_scaling_timescale . Therefore, it struggles to respond effectively to sharp second-level spikes and is infeasible for adjusting hardware at millisecond-level intervals.
In contrast, LLM serving systems operate on a continuous batching basis orca , where requests are scheduled at the granularity of iteration steps usually every tens to hundreds of milliseconds. As the batch size of each step directly impacts token-level latency, variability in request rates must therefore be evaluated at the same short-term timescale. Therefore, high variability within second-level time scales can cause substantial fluctuations in TTFT and TPOT, underscoring the need for serving strategies that can operate effectively at finer timescales, independent of slower hardware scaling mechanisms.
Figure 1: (a) Request rates fluctuations observed on May 10, 2024 from Azure LLM inference trace. (b) Comparison of p90 TPOT across different model precisions, using the Azure LLM inference trace recorded on May 10, 2024.
3.2 Our Proposal: Dual-Precision LLM Serving
To address short-term fluctuations in resource demand during LLM serving, we propose a simple yet effective dual-precision strategy that dynamically switches between FP16 and FP8 execution. When system load is low, FP16 is used to preserve maximum model quality. Under high load conditions, the system falls back to FP8 execution to reduce computational and memory overhead, thereby maintaining responsiveness. Although FP8 may incur slight accuracy degradation, it is preferable to violating SLOs during periods of request surges.
Figure 1 shows p90 TPOT under three precision modes—FP16, FP8, and our dual-precision scheme—using the Llama 3.1 8B model with vLLM on an H100 GPU. Request patterns are based on Microsoft Azure’s LLM inference trace azure_llm_inference_trace , downscaled to 20%. During the 60-second interval with fluctuating load (1–11 req/s, avg. 5.05 req/s), FP16 incurs 19 seconds of SLO violation (TPOT > 33 ms), while FP8 reduces this to 8 seconds. Our dual-precision strategy maintains the same SLO compliance as FP8 while preserving FP16 accuracy for over 68% of the time. This demonstrates its effectiveness in improving latency robustness without sacrificing model quality.
3.3 How to Implement Dual-Precision LLM Serving?
To enable adaptive use of both FP8 and FP16 precision in LLM serving, two straightforward implementation strategies can be considered: separate storage and on-the-fly dequantization.
Separate Storage. One simple approach is to maintain both FP8 and FP16 models in memory and switch between them based on task requirements. While this enables the on-demand selection between lower- and higher-precision models, it results in a 50% memory overhead—an impractical burden for LLM serving systems already constrained by model size. This memory capacity overhead not only limits the maximum model size that can be deployed but also reduces the available memory for key-value (KV) caching, thereby restricting the number of concurrent requests and degrading overall serving throughput.
On-the-fly Quantization. An alternative approach is to store only the FP16 model and perform on-the-fly quantization to enable execution on FP8 tensor cores. Unlike the separate storage method, this approach avoids additional memory capacity overhead. However, it can lead to severely suboptimal GEMM performance in FP8 mode. Despite executing in 8-bit precision, the full 16-bit weights must still be fetched from memory, effectively doubling memory traffic compared to native FP8 execution. As a result, memory bandwidth becomes a bottleneck, limiting the potential throughput gains of FP8 and negating its advantages over the FP16 baseline.
These limitations highlight the need for a more sophisticated dual-precision serving mechanism—one that eliminates additional memory overhead while enabling efficient GEMM execution in both FP16 and FP8 modes.
4 NestedFP
4.1 NestedFP Overview
For efficient dual-precision LLM serving, we propose NestedFP, a scheme that stores only 16-bit weights in memory and extracts 8-bit upper tensor from them when operating in FP8 mode. This approach not only avoids additional memory usage, but also prevents memory traffic amplification—such as fetching 16-bit weights when performing 8-bit GEMM. Figure 2 illustrates the core concept of NestedFP, which addresses the following two key challenges to make this approach viable.
First, the extracted 8-bit representation must maintain accuracy comparable to existing FP8 quantization methods. Naively truncating the upper 8 bits of FP16 values (which use the E5M10 format) produces a low-precision representation similar to E5M2, which offers limited precision compared to the commonly preferred E4M3 format for LLM inference. Note that FP16 is basically E5M10. Moreover, this naive truncation often performs worse due to indiscriminate rounding, further degrading accuracy. This motivates the need for a quantization-aware format that simultaneously supports both FP16 and FP8 execution. Section 4.2 elaborates on this data format design.
Second, kernel efficiency must be maintained across both precisions. Achieving this requires storing the two 8-bit halves of each 16-bit weight as separate tensors. On H100 GPUs, where DRAM is accessed in 64-byte units, extracting individual bytes from packed 16-bit data results in wasted bandwidth. Separating the 8-bit components enables efficient FP8 execution using a single tensor, while FP16 models require reconstructing the original 16-bit weights via a custom GEMM kernel that dynamically fuses the components at runtime. Section 4.3 details this kernel design.
Figure 2: FP16 weight is first pre-processed into 8-bit Upper and Lower parts. Both parts are used for FP16 model execution while only Upper part is used for FP8 model execution.
4.2 Compact yet Effective Data Format for Dual-Precision LLM
Opportunities: Low-Entropy Exponent Bit in FP16 Models. Figure 3 shows the the weight distributions of FP16 linear layers—including QKVO projections and MLPs—across four different LLMs. The vast majority of weight values have absolute magnitudes less than or equal to 1.75, which means that their most significant exponent bits in FP16 are zero. Specifically, as shown in Figure 3, three of the four evaluated models exhibit this trait across all layers. The sole exception, Phi-4, contains some layers that deviate from this pattern, but these account for only 8.75% of its layers. This strong tendency suggests an opportunity for lossless exponent mapping from FP16 to the FP8 E4M3 format, which also uses 4 exponent bits.
Min Max NestedFP Applicable Layers
Llama 3.1 (8B) -0.92 0.91 128 of 128 (100%)
Mistral Nemo Base (12B) -0.92 0.95 160 of 160 (100%)
Phi-4 (14B) -3.60 3.89 146 of 160 (91.25%)
Mistral Small Base (24B) -0.49 0.45 160 of 160 (100%)
Figure 3: (a) Weight distributions of linear layers across different models. (b) Per-model weight range (min/max) and count of linear layers eligible for NestedFP quantization.
Our Proposal. Leveraging the unused exponent bit in FP16, NestedFP restructures FP16 values so that high-quality E4M3 representations can be directly extracted. Figure 4a illustrates the refactoring process. The lower 8 bits of new format of NestedFP is simply constructed by inheriting lower 8 bits of the FP16 mantissa. The upper 8 bits are formed from the sign bit and 4 exponent bits (excluding the most significant bit of the FP16 exponent) of the original FP16 value, along with a 3-bit mantissa obtained by quantizing its 10-bit mantissa.
This simple mapping scheme introduces a fixed scaling factor determined by the difference in exponent bias between the FP16 and E4M3 formats, which is $2^{8}$. In other words, NestedFP effectively applies E4M3 quantization with a scaling factor of $2^{8}$. To quantize the 10-bit FP16 mantissa down to 3 bits, we employ a round-to-nearest-even strategy. Specifically, we examine the 7 least significant bits of the FP16 mantissa (i.e., bits M[4:10]) to determine whether the value lies above or below the midpoint (64). If the value is greater than 64, we round up. If the value equals 64, we round up only when the least significant bit of the resulting 3-bit mantissa (M3) is 1. If the value is less than 64, we round down. This division process is done offline prior to model execution.
FP8 GEMM can be performed directly using the upper 8 bits, whereas FP16 GEMM requires online reconstruction of 16-bit values, introducing additional complexity. Figure 4b illustrates this reconstruction process. Essentially, it concatenates the upper 8 bits and lower 8 bits, but not by simple concatenation. The earlier rounding decision must be reversed by checking an implicit checksum; if the LSB of the upper 8 bits differs from the MSB of the lower 8 bits, a rounding-down correction is applied before recombining the exponent and mantissa bits into a valid FP16 value. This process ensures precise, lossless recovery of the original FP16 values.
(a) Refactoring to NestedFP
(b) Reconstruction of FP16
Figure 4: (a) Pre-processing FP16 weight value into two 8-bit NestedFP components (Offline). (b) On-the-fly reconstruction of FP16 weight values for tensor core execution (Online).
Handling Exception Layers. Since most models do not contain weights with absolute values exceeding 1.75, we leave these values in their original FP16 format without pre-processing. As a result, the corresponding layers are always executed in FP16, consistent with standard FP16 model execution. As shown in Sections 5, this exception handling preserves both model accuracy and end-to-end performance, as these layers constitute only a small portion of the total linear layers.
4.3 GPU Kernel Design
While executing FP8 GEMM in NestedFP is straightforward—simply performing GEMM on the NestedFP upper tensor, FP16 GEMM is more complex. Figure 6 illustrates the FP16 GEMM process in both a standard FP16 CUTLASS kernel and the NestedFP16 kernel. In the standard FP16 kernel, weights are loaded into shared memory, transferred to registers, and directly consumed by tensor cores without any transformation. In contrast, the FP16 path in NestedFP requires a reconstruction step before tensor core execution; weights are stored as two 8-bit tensors, which must be loaded into shared memory, copied to registers, and then reconstructed into full FP16 values through SIMT-level bitwise operations. Only after this reconstruction can the data be used by the tensor cores.
This additional step introduces a source of overhead absent in traditional FP16 pipelines. To prevent it from becoming a performance bottleneck, SIMT operations must be highly efficient, and their latency must be effectively hidden within the broader compute pipeline. We present two key optimization strategies to address this challenge: (1) SIMT operation optimization and (2) pipelining and scheduling optimizations to prevent SIMT operations from stalling the tensor core pipeline.
SIMT Reconstruction Operation Optimization. To mitigate the overhead introduced by SIMT operations, we fuse four 8-bit bitwise operations into a single 32-bit operation, following a strategy similar to any_precision_llm ; fp6_llm . As shown in Figure 6, the reconstruction process combines two 8-bit tensors, $W_{1}$ and $W_{2}$, stored in registers, into a full 16-bit FP16 weight. It begins with a correction step for $W_{1}$, followed by dropped $E_{1}$ restoration, and ends with the final fusion of $W_{1}$ and $W_{2}$. In this scheme, the lower 8 bits of the FP16 mantissa, $M[3:10]$, are taken directly from $W_{2}$. The upper tensor, $W_{1}$, contains $E[2:5]$ and $M[1:2]$, which may require correction depending on rounding. Rounding is detected using a checksum condition: the LSB of $W_{1}$ ($M_{3}^{\prime}$) and the MSB of $W_{2}$ ($M_{3}$) should match. If $M_{3}^{\prime}=1$ and $M_{3}=0$, no correction is needed for $E[2:5]$ and $M[1:2]$. If $M_{3}^{\prime}=0$ and $M_{3}=1$, rounding may have corrupted $E[2:5]$ and $M[1:2]$, and correction is applied by subtracting 1 at the $M_{3}$ bit position. To unify both cases and eliminate branching, we apply a branch-free correction; by subtracting $M_{3}$ from $W_{1}$ and taking only bits that corresponds to bit position $E[2:5]$ and $M[1:2]$ from $W_{1}$. Omitted zero is then restored and the full FP16 value is reconstructed using NVIDIA’s __byte_perm instruction.
Figure 5: FP16 execution flow comparison.
Figure 6: Algorithm of SIMT-Efficient Bitwise Reconstruction of FP16 Weights.
Pipelining and Scheduling Optimizations. The standard FP16 GEMM kernel sources weight elements from registers and employs a two-stage pipeline, overlapping shared memory-to-register transfers with tensor core execution. Following prior work that integrates SIMT operations into the tensor core pipeline fp6_llm , we extend this design to a three-stage pipeline that concurrently overlaps register to shared memory transfers, SIMT operations, and tensor core execution. However, simply adopting a three stage pipeline is insufficient to fully hide the overhead of SIMT operations. Through empirical analysis, we find that distinct optimization strategies are required for non-cooperative and cooperative kernels. In the non-cooperative kernel, four warps jointly execute a single warp group MMA (Matrix Multiply Accumulate), whereas the cooperative kernel employs eight warps to execute two warp group MMAs concurrently.
While shared memory transfers and tensor core executions are asynchronous operations—allowing subsequent instructions to be issued without waiting for completion—SIMT operations are synchronous and block instruction issue during their execution. In the non-cooperative kernel, we advance the scheduling of shared memory to register transfers relative to the naive three stage pipeline and issue multiple copies in bulk. This approach not only allows multiple shared memory to register copy concurrent with SIMT execution, but also allows SIMT instruction to be issued without delay as operands are pre-loaded into registers. However, in the cooperative kernel, which leverages additional compute parallelism by executing two warp group MMAs concurrently, warp groups contend for shared resources such as SIMT pipelines, tensor core pipelines, and shared memory bandwidth. This contention reduces the effectiveness of the optimizations used in the non-cooperative kernel. To address this, we explored several alternatives and found that inserting a compiler-level fence—specifically, an NVVM code motion barrier—after the transformation stage significantly improves performance. This fence inhibits compiler optimizations that may delay or reorder memory operations and serves as a signal for the imminent use of a transformed input in a register, preventing the compiler from eliminating or deferring its materialization.
5 Evaluation
5.1 NestedFP8 Accuracy
To evaluate the accuracy of our NestedFP8 model, we compare it with an FP8 model using commonly adopted quantization configurations—per-channel weight quantization and per-token activation quantization, both scaled by the absolute maximum value. We evaluate on four models (Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), and Mistral Small(24B)) using three downstream tasks: Minerva Math, MMLU Pro, and BBH. All models are integrated into the vLLM framework, and accuracy is assessed using the LM Evaluation Harness lm-evaluation-harness framework. As noted in Section 4.2, 8.75% of linear layers in Phi-4 are not applicable to NestedFP and are thus executed in FP16. Our method internally applies a single global scaling factor (256) to all weight tensors. For activation quantization, we apply per-tensor scaling based on the absolute maximum value. As shown in Table 2, our method achieves accuracy comparable to the FP8 baseline despite foregoing fine-grained quantization, demonstrating that NestedFP8 can preserve accuracy.
Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.9 16.8 42.9 41.3 34.6 35.7 MMLU Pro 32.8 33.0 34.9 35.0 52.7 53.3 53.6 53.7 BBH 38.6 38.8 41.2 40.6 27.1 27.9 51.8 51.4
Table 2: Accuracy comparison between FP8 baseline and NestedFP8. FP8(B) denotes FP8 baseline, and FP8(N) denotes NestedFP8.
5.2 NestedFP16 Kernel Performance
Comparison over Vanilla FP16 Kernel. To compare the performance of NestedFP16 kernel with normal FP16 GEMM kernel fairly, we implemented a highly optimized FP16 baseline using the same CUTLASS codebase. This baseline selects the best configuration for each GEMM shape through exhaustive search over the design space, closely matching cuBLAS performance (validated in Appendix). Our kernel design space includes non-cooperative and cooperative variants. For non-cooperative kernels, we perform a grid search over tile shapes with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$. Cooperative kernels use larger tiles ($T_{n}\in{128,256}$), and configurations that fail to compile are excluded. Non-cooperative kernels use $(1,1,1)$ clusters, while cooperative ones use $(2,1,1)$, assigning two thread blocks per GPU Processing Cluster. We use the standard data parallel scheduler for non-cooperative kernels and both the standard data parallel and Stream-K streamk schedulers for cooperative ones. All benchmarks are conducted on an NVIDIA H100 SXM GPU.
Our NestedFP16 kernels are optimized using the same design space as the FP16 baseline, ensuring that each GEMM shape is paired with its optimal kernel configuration. We benchmark both kernels across all linear-layer GEMM shapes used in LLaMA 3.1–8B, Mistral Nemo, Phi-4, and Mistral Small, each of which involves four distinct $(N,K)$ shapes. The batch dimension $(M)$, determined by the number of input tokens during inference, is swept from 32 to 2048 in increments of 32, yielding a total of 882 GEMM configurations across 14 unique $(N,K)$ shapes.
Figure 7 compares the performance of the largest $(N,K)$ GEMM configurations across the four models. As shown, our kernel consistently maintains low SIMT overheads across all $M$ values. Compared to the CUTLASS baseline, the average relative overheads across all GEMM configurations of each model are 5.69% (LLaMA 3.1–8B), 6.83% (Mistral Nemo), 6.65% (Phi-4), and 6.71% (Mistral Small). Additional results for other GEMM shapes, as well as comparisons with PyTorch W16A16 kernels, which exhibit similar trends, are provided in Appendix. While certain kernel performance regarding certain $M$ show more variability, the average overhead remains modest, demonstrating the practicality of our approach across diverse shapes in real LLM inference scenario.
Impact of Optimization Techniques. To quantify the impact of successive kernel-level optimizations, we measure latency for a matrix multiplication of shape $M\times 5120\times 32768$. Figure 7 illustrates the latency reduction achieved across three optimization stages. Level 1 introduces a three-stage pipeline that overlaps shared memory transfers with SIMT execution and tensor core operations. Level 2 incorporates SIMT-level FP16 reconstruction optimization. This transition yields a substantial 38.3% reduction in latency relative to Level 1. Level 3 applies additional pipelining and scheduling optimizations resulting in a further 11.0% reduction compared to Level 2. Each stage contributes a measurable improvement, highlighting the cumulative benefits of the proposed optimization strategies.
Figure 7: (a) CUTLASS baseline and NestedFP16 kernel performance comparison. (b) Kernel latency reduction with different levels of optimization strategies.
5.3 End-to-End Performance
We integrated NestedFP into the vLLM framework as a quantization configuration. The system supports per-iteration precision switching, allowing the scheduler to dynamically select FP16 or FP8 based on workload and SLO demands. To ensure fair comparison, both FP16 baselines and NestedFP models were compiled using the vLLM 0.8.3 V1 engine and PyTorch 2.6.0+cu12.4. Experiments were run on a single NVIDIA H100 GPU with fixed request sizes (256 input, 512 output tokens) and batch sizes ($M$) ranging from 32 to 512 to assess throughput scaling. Throughput is reported in tokens per second (tok/s). Figure 8 shows end-to-end performance under FP16 and NestedFP. The FP16 runs use torch.matmul, which dispatches to cuBLAS by default within the vLLM framework. Compared to this baseline, NestedFP16 introduces only minor overhead—4.50% (LLaMA 3.1–8B), 4.51% (Mistral Nemo), 2.69% (Phi-4), and 3.91% (Mistral Small)—generally lower than kernel-level overheads, likely due to amortization across non-GEMM components. In contrast, NestedFP8 achieves clear speedups over NestedFP16: 1.24$\times$, 1.32$\times$, 1.29$\times$, and 1.53$\times$, respectively. Larger models gain more, reflecting the growing cost of matrix multiplication. Overall, NestedFP improves efficiency for large-scale LLMs.
Figure 8: End-to-end performance comparison with four models using FP16, NestedFP16 and NestedFP8.
6 Conclusion
In this paper, we introduce NestedFP, a precision-aware serving framework designed to better satisfy service-level objectives (SLOs) under dynamic request conditions. We propose a novel approach for decomposing FP16 weights into two 8-bit tensors, enabling one to be used directly as an FP8 representation. Our FP8 format achieves accuracy comparable to that of standard quantized FP8 models. In addition, we develop a custom 16-bit GEMM kernel tailored to this dual-format representation. Experimental results show that our kernel achieves an average performance gap of 6.47% compared to the CUTLASS baseline kernel. Furthermore, by integrating our method into the vLLM framework, we observe up to 1.55$\times$ speedup in FP8 model execution relative to FP16 model execution, while incurring average 3.90% overhead across models when executing in FP16 mode. These results demonstrate that NestedFP successfully bridges the gap between FP16 accuracy and FP8 efficiency, requiring the same memory footprint as a standard FP16 model while supporting both precision modes. By enabling iteration-level precision selection, our approach allows production LLM serving systems to dynamically balance performance and accuracy requirements under varying workload conditions, providing a flexible and efficient solution for real-world deployment scenarios.
Appendix A Evaluation Details
We conducted our evaluation using the Mistral Nemo Base 2407 and Mistral Small 24B Base 2501 models. For accuracy assessment, we employed the Math Verify Metric on the Minerva Math task, the Open LLM Leaderboard configuration (5-shot, multiple choice) for the MMLU Pro task, and a zero-shot evaluation setting for the BBH task. Kernel performance was measured using NVIDIA Nsight Compute. In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$ as it provides more robust performance. During performance measurement, we flushed all caches using the cache-control option, but did not fix the GPU clock frequency using the clock-control option.
Appendix B Extended Kernel Performance for NestedFP16
Figure 9: Performance comparison between our CUTLASS baseline and NestedFP16 kernel across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
Figure 9 compares our method against the CUTLASS FP16 baseline across all 14 GEMM shapes for the four evaluated models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small. The matrix dimension $M$ is varied in increments of 32, ranging from 32 to 2048. Our results show that NestedFP consistently incurs only moderate overhead across all GEMM configurations, with an average performance difference of 6.1%. The per-graph average overhead ranges from 4.3% to 7.2%, demonstrating stable performance regardless of matrix shape. This consistency highlights the effectiveness of our kernel-level optimizations.
Appendix C Extended End-to-End Throughput Evaluation
(a) (32, 512)
(b) (1024, 512)
(c) (32, 32)
(d) (1024, 32)
Figure 10: Extended end-to-end throughput evaluation across different input/output token configurations. Each subplot shows throughput comparison for FP16, NestedFP16, FP8 and NestedFP8 across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
We extend our end-to-end evaluation to further demonstrate the robustness of NestedFP. Figure 10 presents results for four additional input/output configurations: (32, 512), (1024, 512), (32, 32), and (1024, 32). Across all settings, NestedFP16 exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 3.88% and a maximum of 8.22%. On a per-model basis, the average overhead is 5.15% for LLaMA 3.1 (8B), 4.19% for Mistral-Nemo (12B), 3.35% for Mistral-Small (24B), and 2.81% for Phi-4 (14B). In contrast, NestedFP8 achieves significant speedups over NestedFP16, with average throughput gains of 1.26$\times$ for LLaMA 3.1, 1.33$\times$ for Mistral-Nemo, 1.53$\times$ for Mistral-Small, and 1.36$\times$ for Phi-4 across all input/output shapes. Notably, the largest model, Mistral-Small, benefits the most from the NestedFP8 quantization. These results indicate that the performance gains of NestedFP are consistent and robust across a wide range of input/output configurations.
When compared to Torch FP8, NestedFP8 maintains competitive performance, achieving 96.84% of Torch FP8’s throughput for LLaMA 3.1, 97.76% for Mistral-Nemo, 98.79% for Mistral-Small, and 96.77% for Phi-4. Our limited search space shows potential gain that can be achieved with further FP8 kernel optimization. We note that we did adopt Stream-K scheduling for NestedFP8.
Our evaluation setup mirrors the configuration used in the main throughput experiments. All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.3 V1 engine with PyTorch 2.6.0+cu12.4 for both the baseline and our proposed NestedFP implementations. Chunked prefill, enabled by default in the vLLM V1 engine, was used throughout the evaluation. This feature segments the prefill phase into smaller chunks, which are processed concurrently with decoding tokens within the same batch. At each iteration, the GEMM operation shape $(M,N,K)$ is determined based on the number of tokens grouped in that step, with $M$ denoting the total token count. To systematically evaluate throughput under different batch sizes, we adjusted the max_num_batched_token and max_num_seqs parameters to constrain the batch size accordingly.
Appendix D CUTLASS FP16 GEMM Kernel
We extend the open-source NVIDIA CUTLASS 3.6 library to support our custom format. We begin by describing the efficient design of GEMM kernels within CUTLASS for the NVIDIA H100 GPU and introduce key components that define our design space. In the following section, we benchmark our CUTLASS kernels against PyTorch to validate strong performance of CUTLASS baseline.
D.1 GEMM Kernel Design
Figure 11: (a) Data parallel scheduling. A single output tile is processed by a single thread block. (b) Stream-K scheduling. A single output tile is processed collaboratively across multiple thread blocks. (c) DRAM to shared memory copy handled by 4 warps in producer warp group and TMA engine.
Figure 12: (a) Non-cooperative kernel. A single warp group is responsible for processing entire input tile. (b) Cooperative kernel. Two warp groups are each responsible for processing half of the first input tile and corresponding output tile.
GEMM operates on a weight matrix of shape $N\times K$ and an activation matrix of shape $M\times K$. To parallelize this computation, CUTLASS partitions the input matrices into tiles of size $T_{n}\times T_{k}$ and $T_{m}\times T_{k}$, respectively. These tiles define the granularity at which work is distributed across thread blocks.
As illustrated in Figure 11, each input tile is processed along the $K$ dimension to produce a corresponding output tile, which is accumulated in registers. CUTLASS employs a persistent tile scheduler, assigning each thread block to a fixed streaming multiprocessor (SM) for the duration of execution. In data-parallel scheduling, each SM processes the full $K$ dimension for a given output tile independently. In contrast, stream-K scheduling enables multiple thread blocks to collaboratively compute a single output tile by dividing the $K$-dimension workload across thread blocks, as shown in Figure 11. The partial results are later merged to form the final output. In all cases, the computed output tiles are written back to global memory (DRAM).
To hide DRAM latency and enable compute–memory overlap, tiles are asynchronously loaded from global memory into shared memory. As shown in Figure 11, shared memory is partitioned to hold multiple tiles, allowing future iterations to be staged while the current tile is being processed. On NVIDIA H100 GPUs, this transfer is managed by the Tensor Memory Accelerator (TMA) engine. CUTLASS uses warp specialization to pipeline these operations: one warp group (four warps) handles TMA data transfers (producers), while one or two other warp groups (consumers) execute matrix-multiply-accumulate (MMA) operations, depending on the kernel configuration.
Once data is available in shared memory, the kernel proceeds with MMA using one of two configurations. In the RS (register-shared) kernel, weights are loaded into registers and processed with activations from shared memory—this approach enables bit-level manipulations required by our custom format. In contrast, SS (shared-shared) kernels source both operands directly from shared memory. On Hopper, SM90 MMA instructions require the second operand to come from shared memory. MMA operations are executed at the warp group granularity, requiring all warps in a group to be ready before issuing the instruction. Consequently, SIMT-level manipulations on weights must complete across all four warps before any warp group MMA operation begins.
CUTLASS supports both non-cooperative and cooperative kernel modes. Non-cooperative kernels use a single consumer warp group, while cooperative kernels employ two consumer groups that share activations but process distinct weight tiles. The behavioral differences between these modes are illustrated in Figure 12.
D.2 CUTLASS FP16 Baseline and cuBLAS W16A16 Kernel Comparison
Figure 13: Performance comparison between PyTorch torch.matmul and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes. Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
To validate the robustness of our CUTLASS baseline, we compare its performance against the cuBLAS kernels employed by PyTorch 2.6.0 with CUDA 12.4. Using a weak baseline could underestimate the overhead introduced by NestedFP, leading to overly optimistic conclusions about its practical efficiency. Figure 13 provides a detailed comparison across 14 $(N,K)$ configurations, extracted from the linear layers of the four models evaluated in our main experiments—Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small. For each configuration, we sweep the batch size $M$ from 32 to 2048 in increments of 32, covering the typical range encountered in LLM inference workloads.
Our results show that the FP16 baseline closely matches the performance of the cuBLAS implementation, with an average performance difference of just 1.8% across all 14 GEMM shapes. In 11 of the 14 configurations, our baseline achieves higher average performance than PyTorch, with performance differences reaching up to 3.3%. For the remaining three shapes—corresponding to smaller GEMMs—PyTorch slightly outperforms our implementation, with differences ranging from 0.3% to 0.9%, well within the bounds of normal GPU performance variability. These results confirm that our CUTLASS baseline is not only competitive but often exceeds production-level kernel performance, ensuring a reliable and fair comparison when evaluating NestedFP’s overhead.
Appendix E Broad Applicability of NestedFP
Model GEMM1 GEMM2 GEMM3 GEMM4 Total
CodeLlama 7B 96/96 32/32 64/64 31/32 223/224 (99.6%)
CodeLlama 13B 120/120 40/40 80/80 37/40 277/280 (98.9%)
Gemma 3 4B 207/264 64/88 123/176 34/34 429/563 (76.2%)
Gemma 3 12B 249/306 78/102 151/204 48/48 527/661 (79.7%)
Gemma 3 27B 291/348 92/116 179/232 62/62 625/759 (82.3%)
Llama 3.1 8B 96/96 32/32 64/64 32/32 224/224 (100.0%)
Llama 3.1 70B 224/240 80/80 141/160 78/80 523/560 (93.4%)
Mistral Nemo 12B 120/120 40/40 80/80 40/40 280/280 (100.0%)
Mistral Small 24B 120/120 40/40 80/80 40/40 280/280 (100.0%)
Phi-3.5 Mini 26/32 31/32 31/32 24/32 112/128 (87.5%)
Phi-4 14B 40/40 38/40 40/40 28/40 146/160 (91.2%)
Qwen 3 8B 108/108 35/36 72/72 34/36 249/252 (98.8%)
Qwen 3 14B 120/120 40/40 80/80 38/40 278/280 (99.3%)
Qwen 3 32B 192/192 63/64 127/128 56/64 438/448 (97.8%)
Table 3: Layer-wise applicability of NestedFP across models. Format X/Y indicates X applicable layers out of Y total layers for each GEMM type.
Our scheme is broadly applicable to a wide range of LLMs beyond those evaluated in the main experiments. In Table 3, we analyze ten additional models that exhibit the key characteristic: the majority of their weight values have absolute magnitudes less than or equal to 1.75. For each model, we examine four distinct GEMM shapes corresponding to different linear layer types: GEMM1 (QKV projections), GEMM2 (output projections), GEMM3 (MLP gate/up projections), and GEMM4 (MLP down projections). Each entry is shown as X/Y, indicating that X out of Y layers of that type satisfy the weight magnitude constraint (absolute value $\leq$ 1.75).
Our analysis indicates that most models exhibit high applicability rates. The Gemma 3 family shows the lowest rates (76.2%–82.3%), primarily due to multi-modal projection layers containing weights with significantly higher magnitudes (up to 26.25). Among the various GEMM types, GEMM4 (MLP down projections) tends to include more layers exceeding the threshold in models such as Phi-4 14B (70.0% applicability) and Qwen 3 32B (87.5%). While the majority of models have maximum weight values below 3.0, notable outliers include Llama 3.1 70B (maximum value of 93.0) and the Gemma 3 series (maximum value of 26.25). These extreme values are typically confined to a small number of layers and have limited impact on the overall applicability of our method. The consistently high applicability rates across diverse model families affirm the robustness and generalizability of our approach’s core assumption regarding weight distributions.

### v2

@@
 Abstract
-Large Language Models (LLMs) are playing a crucial role in latency-critical, high-throughput services like virtual assistants and code generation.
-While techniques such as continuous batching and paged attention address service-level objectives (SLOs), and quantization methods accelerate inference, the dynamic and efficient adaptation of precision at runtime remains a significant, largely underexplored challenge.
-The emergence of hardware support for FP8 arithmetic, offering up to 2$\times$ the throughput of FP16, presents an attractive opportunity for interactive LLM serving.
-However, current approaches like co-deploying FP8 and FP16 models suffer from increased storage overhead and fail to unlock FP8’s full potential.
-To address these limitations, we introduce NestedFP, a novel precision-adaptive serving technique enabling seamless FP8 and FP16 inference from a single 16-bit model representation, thereby incurring no additional memory cost.
-NestedFP decomposes each FP16 weight into two 8-bit components, facilitating efficient FP8 execution while preserving full FP16 accuracy.
-We demonstrate the practical viability of our approach by implementing a custom CUTLASS-based GEMM kernel that reconstructs FP16 operands on-the-fly, integrated within the vLLM serving framework.
-Our evaluation shows that NestedFP delivers up to 1.55$\times$ throughput improvement in FP8 mode with negligible accuracy degradation compared to FP16 precision, while introducing only 3.9% performance overhead on average in FP16 mode across various models.
-NestedFP thus provides a flexible foundation for dynamic, SLO-aware precision selection, paving the way for more scalable and efficient LLM serving under bursty and heterogeneous workloads.
+Meeting service-level objectives (SLOs) in Large Language Models (LLMs) serving is critical, but managing the high variability in load presents a significant challenge.
+Recent advancements in FP8 inference, backed by native hardware support, offer a potential solution: executing FP16 models by default, while switching to FP8 models during sudden load surges to achieve higher throughput at the cost of a slight quality degradation.
+Although this approach facilitates effective SLO management, it introduces additional memory overhead due to storing two versions of the same model.
+In response, this paper proposes NestedFP, an LLM serving technique that supports both FP16 and FP8 models in a memory-efficient manner by overlaying FP8 parameters onto FP16 parameters, allowing both models to share the same FP16 memory footprint.
+By leveraging a compact data format for the overlay and a specialized GEMM kernel optimized for this format, NestedFP ensures minimal degradation in both model quality and inference throughput across both FP8 and FP16 modes.
+NestedFP provides a flexible platform for dynamic, SLO-aware precision selection.
+The code is available at https://github.com/SNU-ARC/NestedFP.
 1 Introduction
-Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information.
+Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information gpt4_technical_report ; gpt3_technical_report ; llama3_technical_report .
 Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand.
-For example, OpenAI reportedly generates over a billion words per day—equivalent to more than one million words per second marlin —with production query rates frequently surpassing 200 requests per second scale_llm .
-This widespread deployment in latency-sensitive, user-facing services has intensified the need for inference infrastructures that are not only scalable and efficient, but also capable of satisfying stringent latency and throughput constraints.
-The latency of LLM inference is commonly evaluated using two standard metrics: Time To First Token (TTFT), which measures the prefill latency, and Time Per Output Token (TPOT), which captures the average latency for generating each output token.
-For interactive applications, industry-standard service-level objectives (SLOs) typically require TTFT to be under 200 ms and TPOT to remain below 33.3 ms in order to ensure a responsive user experience baseten_slo .
-To meet these stringent latency targets under growing demand, several optimization techniques have been proposed.
-ORCA orca introduces efficient batching mechanisms that consolidate linear layer execution across both the prefill and decode phases.
-PagedAttention paged_attention addresses memory inefficiencies arising from unpredictable output sequence lengths by reducing over-allocation, thereby increasing the number of concurrently served requests.
-More recently, Sarathi-Serve sarathi_serve proposes chunked execution of the prefill phase and constructs hybrid batches combining prefill and decode requests to further improve scheduling efficiency.
-Recent studies have shown that the FP8 format can achieve accuracy comparable to FP16 across a variety of large language model tasks fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 .
-The introduction of native FP8 support on modern accelerators—such as NVIDIA Hopper GPUs and Intel Gaudi HPUs—has made FP8 increasingly practical for large-scale inference.
-On NVIDIA Hopper, FP8-compatible tensor cores can offer up to 2× higher peak throughput compared to their FP16 counterparts, making FP8 particularly advantageous for latency-sensitive or bursty workloads.
-In this work, we propose leveraging both FP8 and FP16 representations to better satisfy service-level objectives (SLOs).
-However, co-deploying models in both formats is non-trivial, as it is difficult to ensure robust accuracy and performance without introducing additional storage overhead beyond the originally required 16-bit model.
-We present NestedFP, a technique that enables dynamic selection between FP8 and FP16 precision from a single, unified model representation.
-By analyzing the value distribution of FP16 weights, our approach restructures each weight tensor into two 8-bit components, enabling accurate FP16 reconstruction while supporting efficient FP8 execution—all with no additional memory storage.
-To realize this format, we develop a custom General Matrix Multiplication (GEMM) kernel built on the CUTLASS library cutlass , which reconstructs FP16 values from the two 8-bit components on-the-fly during GEMM kernel execution.
-The kernel incurs an average performance overhead of 6.47% across varying GEMM shapes, relative to the vanilla CUTLASS kernels, respectively.
-Moreover, our 8-bit tensor for FP8 inference achieves FP8 inference accuracy comparable to that of standard FP8 quantized models.
-We integrate our custom kernel into the vLLM serving framework vllm and show that it incurs only 3.90% end-to-end performance overhead on average across models in FP16 mode, while delivering up to 1.55$\times$ performance improvements when executing in FP8 mode.
-NestedFP provides a flexible and memory-efficient foundation for adaptive precision serving, making it well-suited for fluctuating request rates and stringent latency constraints.
+For example, OpenAI’s production-scale models process roughly 100 billion words per day openai_billion_token and production query rates frequently surpass 200 requests per second chatgpt_fact ; scale_llm .
+To accommodate such large-scale applications, research efforts have focused not only on improving serving throughput orca ; paged_attention but also on meeting service-level objectives (SLOs) sarathi_serve ; llumnix ; sola_llm .
+One of the key challenges in achieving SLOs for LLM serving lies in handling dynamic load.
+Request rates, as well as input and output sequence lengths, can fluctuate significantly over short time intervals.
+Simply overprovisioning compute resources may mitigate this issue but results in substantial system cost inefficiency.
+As a promising alternative, we explore an approach we call dual-precision LLM serving, which utilizes FP8 together with FP16, the de facto standard number format for weight parameters.
+While FP8 may incur a slight degradation in model quality, it offers up to 2× higher peak throughput compared to FP16 fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 .
+By operating in FP16 mode under normal conditions and switching to FP8 mode only during sudden load surges, we can effectively manage dynamic workloads without resource overprovisioning.
+This approach achieves a better balance between SLO attainment and service quality than relying solely on either FP16 or FP8.
+Co-deploying models in both formats, however, is non-trivial.
+Simply keeping both in memory is infeasible due to excessive memory capacity overhead.
+Storing only FP16 weights and quantizing them to FP8 on the fly when needed results in highly suboptimal FP8 throughput.
+Therefore, an effective solution is needed to support dual-precision LLM serving without incurring memory overhead or throughput degradation.
+In this paper, we present NestedFP, a technique that enables efficient dynamic selection between FP8 and FP16 precision from a single unified model representation.
+NestedFP decomposes each 16-bit weight tensor into two 8-bit components, allowing FP8 inference without additional memory overhead or throughput degradation.
+Since FP16 mode requires reconstructing 16-bit values by merging the two 8-bit components in this approach, we develop a custom General Matrix Multiplication (GEMM) kernel, built on the CUTLASS library cutlass , that performs accurate on-the-fly FP16 reconstruction during kernel execution.
+According to our evaluation, NestedFP enables high-quality, dual-precision LLM serving efficiently: FP8 models generated from decomposed FP16 weights match the quality of existing FP8 quantization methods, and our FP16 GEMM kernel incurs only a 6.47% average performance gap relative to the CUTLASS baseline, which translates to 4.98% end-to-end overhead across models.
+Our contributions are summarized as follows:
+  • We propose dual-precision LLM serving as an effective strategy to handle dynamic load fluctuations.
+  • We introduce a novel data format that supports both high-quality FP16 and FP8 inference in a memory-efficient manner.
+  • We develop a custom FP16 GEMM kernel for this data format, which performs on-the-fly FP16 value reconstruction during execution.
+  • We demonstrate that NestedFP enables high-quality dual-precision LLM serving with no memory overhead and high throughput, significantly improving SLO attainment compared to FP16-only baselines.
 2 Background
 2.1 Floating Point Representation
-[…]
+A floating point format with $x$ exponent bits and $y$ mantissa bits is denoted as ExxMyy.
+The exponent width $x$ determines the dynamic range of representable values, while the mantissa width $y$ governs numerical precision.
+A floating point value $X_{FP}$ in this format can be expressed as shown in Equation 1, where $S$ is the sign bit, $M_{j}$ are the mantissa bits, $E$ is the unsigned integer formed by the exponent bits, and $b$ is the exponent bias.
 $X_{FP}=(-1)^{S}\cdot\left(1+\sum_{j=1}^{y}M_{j}\cdot 2^{-j}\right)\cdot 2^{E-b},\quad E\in\{0,1,\ldots,2^{x}-1\}$ (1)
@@
 Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs investigation_fp8_across_accelerators .
-Leveraging this support, FP8 GEMM operations can achieve up to 2×2$\times$2 × speedup compared to their FP16 counterparts.
+Leveraging this support, FP8 GEMM operations can achieve up to $2\times$ speedup compared to their FP16 counterparts.
 This growing availability is accelerating the adoption of floating-point quantization—particularly FP8—further highlighting its relevance for efficient LLM deployment.
@@
 FP8 Results.
-While FP8 quantization offers significant efficiency benefits and shows promising results, it still lags behind FP16 in terms of accuracy.
-Table 1 presents a comparison between FP8 and FP16 across various benchmarks that require sophisticated reasoning capabilities.
-The FP8 results are obtained using the vLLM framework vllm , with E4M3 quantization applied using per-channel weight scaling and per-token activation scaling based on absolute maximum value.
-Quantization is applied exclusively to linear layers, with embedding layers left in higher precision.
-FP8 generally exhibits a measurable drop in accuracy, highlighting its current limitations for tasks that demand high output fidelity.
-Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ Minerva Math 18.1 17.6 --0.54 17.1 16.9 --0.20 42.7 42.9 $+$0.16 35.4 34.6 --0.84 MMLU Pro 33.3 32.8 --0.49 35.4 34.9 --0.44 53.1 52.7 --0.37 54.3 53.6 --0.71 BBH 39.5 38.6 --0.83 41.7 41.2 --0.49 27.5 27.1 --0.38 52.5 51.8 --0.65
+While FP8 inference offers significant throughput gains, it comes with the potential caveat of quality degradation due to the inevitable loss of information associated with reduced precision dfloat11 .
+Table 1 presents a comparison between FP8 and FP16 across various models and benchmarks.
+Specifically, for FP8 quantization, we use the E4M3 format with per-token activation and per-channel weight quantization for all linear layers, using the vLLM framework vllm .
+A consistent accuracy degradation is observed in FP8 models, which clearly shows the trade-off between throughput and model quality in FP8 quantization.
+Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ FP16 FP8 $\Delta$ Minerva Math 18.2 17.6 -0.60 17.2 16.5 -0.72 42.7 42.9 +0.14 35.4 34.2 -1.14 MMLU Pro 33.3 32.8 -0.49 35.4 35.3 -0.06 53.1 52.7 -0.36 54.3 53.6 -0.77 BBH 39.5 38.6 -0.86 41.6 40.6 -0.97 27.5 27.1 -0.40 52.4 50.8 -1.62
 Table 1: Downstream task accuracy for FP8 quantized models relative to FP16 models.
 3 Motivation
-3.1 Challenge: Dynamic Resource Demand in LLM Serving
+3.1 Challenge: Dynamic Load in LLM Serving
 LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests andes ; hygen .
-These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in TTFT and TPOT.
-Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet service-level objectives (SLOs).
-To quantify the extent of this variability, we analyzed production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace .
-These traces contain timestamps of incoming requests, from which we constructed the per-second request rate on May 10, 2024, as shown in Figure 1.
-Our analysis reveals that the request rate ranged from 0 to 100 requests/s across the entire day, and exhibited intra-day variations of up to 5.8-fold within the most variable 1-hour window (14:00–15:00 UTC; min: 17, max: 98) and 3.2-fold within the most variable 1-minute window (18:12–18:13 UTC; min: 31, max: 98).
-To cope with such highly variable request patterns, auto scaling is a commonly used to automatically adjusts computational resources based on real-time demand auto_scaling .
-However, auto scaling is feasible mainly in cloud environments with sufficient available resources, and is typically limited to time scales of few minutes to hours auto_scaling_timescale .
-Therefore, it struggles to respond effectively to sharp second-level spikes and is infeasible for adjusting hardware at millisecond-level intervals.
-In contrast, LLM serving systems operate on a continuous batching basis orca , where requests are scheduled at the granularity of iteration steps usually every tens to hundreds of milliseconds.
-As the batch size of each step directly impacts token-level latency, variability in request rates must therefore be evaluated at the same short-term timescale.
-Therefore, high variability within second-level time scales can cause substantial fluctuations in TTFT and TPOT, underscoring the need for serving strategies that can operate effectively at finer timescales, independent of slower hardware scaling mechanisms.
-Figure 1: (a) Request rates fluctuations observed on May 10, 2024 from Azure LLM inference trace.
+These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in Time To First Token (TTFT) and Time Per Output Token (TPOT).
+Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet SLOs.
+To quantify the extent of this variability, we analyze production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace .
+Specifically, we measure how the request rate fluctuates over a one-hour interval (00:00–01:00 UTC, May 10, 2024, in the trace).
+Figure 1 presents the results.
+The request rate exhibits nearly a fivefold variation between the lowest and highest load periods.
+Moreover, as shown in the one-minute zoom-in view, these fluctuations occur at high temporal frequency, indicating second-level variability.
+A classic approach to handling such load variations is auto-scaling, which adaptively adjusts computational resources (e.g., increasing or decreasing the number of GPUs in a cluster) auto_scaling .
+However, auto-scaling is primarily effective in cloud environments with abundant resources and typically operates on minute-to-hour timescales auto_scaling_timescale .
+Consequently, it is inadequate for LLM serving systems, which must respond to sharp, second-level load spikes and scale throughput within millisecond-level intervals to meet SLO requirements.
+Figure 1: (a) Request rate fluctuations during the first hour of May 10, 2024, from the Azure LLM inference trace.
+The inset plot highlights 1-minute variability from the 30-minute mark.
 (b) Comparison of p90 TPOT across different model precisions, using the Azure LLM inference trace recorded on May 10, 2024.
 3.2 Our Proposal: Dual-Precision LLM Serving
-To address short-term fluctuations in resource demand during LLM serving, we propose a simple yet effective dual-precision strategy that dynamically switches between FP16 and FP8 execution.
-When system load is low, FP16 is used to preserve maximum model quality.
-Under high load conditions, the system falls back to FP8 execution to reduce computational and memory overhead, thereby maintaining responsiveness.
-Although FP8 may incur slight accuracy degradation, it is preferable to violating SLOs during periods of request surges.
-Figure 1 shows p90 TPOT under three precision modes—FP16, FP8, and our dual-precision scheme—using the Llama 3.1 8B model with vLLM on an H100 GPU.
-Request patterns are based on Microsoft Azure’s LLM inference trace azure_llm_inference_trace , downscaled to 20%.
-During the 60-second interval with fluctuating load (1–11 req/s, avg.
-5.05 req/s), FP16 incurs 19 seconds of SLO violation (TPOT > 33 ms), while FP8 reduces this to 8 seconds.
-Our dual-precision strategy maintains the same SLO compliance as FP8 while preserving FP16 accuracy for over 68% of the time.
-This demonstrates its effectiveness in improving latency robustness without sacrificing model quality.
+To address short-term fluctuations in resource demand during LLM serving, we propose a simple yet effective dual-precision strategy that dynamically switches between FP16 and FP8 modes.
+When the system load is low, FP16 is used to preserve maximum model quality.
+Under high load conditions, the system falls back to FP8 mode to increase throughput, thereby maintaining responsiveness.
+Although FP8 mode may incur slight accuracy degradation, it is often preferable to violating SLOs during periods of request surges.
+Figure 1 presents the per-second p90 TPOT under three precision schemes: FP16, FP8, and the proposed dual-precision format, using the Llama 3.1 8B model on an H100 GPU with vLLM.
+The request patterns are extracted from a 60-second segment of Microsoft Azure’s LLM inference trace azure_llm_inference_trace , scaled down to 20% of the original load.
+The TPOT SLO threshold is set to 33 ms baseten_slo .
+As shown, the FP16 model exhibits more frequent TPOT spikes that exceed the TPOT SLO threshold compared to the FP8 model.
+To further quantify the difference, we analyze the rate of SLO violations across the same trace: 35.8% of requests under FP16 exceeded the threshold, whereas FP8 reduced this to 17.3%.
+In short, the proposed dual-precision inference can achieve FP8-level SLO compliance while minimizing quality degradation.
 3.3 How to Implement Dual-Precision LLM Serving?
-To enable adaptive use of both FP8 and FP16 precision in LLM serving, two straightforward implementation strategies can be considered: separate storage and on-the-fly dequantization.
+To enable dynamic switching between FP16 and FP8 modes in LLM serving, two straightforward implementation strategies can be considered: separate storage and on-the-fly dequantization.
 Separate Storage.
-One simple approach is to maintain both FP8 and FP16 models in memory and switch between them based on task requirements.
-While this enables the on-demand selection between lower- and higher-precision models, it results in a 50% memory overhead—an impractical burden for LLM serving systems already constrained by model size.
-This memory capacity overhead not only limits the maximum model size that can be deployed but also reduces the available memory for key-value (KV) caching, thereby restricting the number of concurrent requests and degrading overall serving throughput.
+One simple approach is to maintain both FP16 and FP8 models in memory and switch between them according to system load fluctuations.
+While this enables fast switching, it introduces a 50% memory overhead—a significant burden for LLM serving systems already constrained by model size.
+This memory capacity overhead not only limits the maximum deployable model size but also reduces the available memory for key-value (KV) caching, thereby restricting the number of concurrent requests and ultimately degrading overall serving throughput.
 On-the-fly Quantization.
-An alternative approach is to store only the FP16 model and perform on-the-fly quantization to enable execution on FP8 tensor cores.
-Unlike the separate storage method, this approach avoids additional memory capacity overhead.
-However, it can lead to severely suboptimal GEMM performance in FP8 mode.
-Despite executing in 8-bit precision, the full 16-bit weights must still be fetched from memory, effectively doubling memory traffic compared to native FP8 execution.
-As a result, memory bandwidth becomes a bottleneck, limiting the potential throughput gains of FP8 and negating its advantages over the FP16 baseline.
-These limitations highlight the need for a more sophisticated dual-precision serving mechanism—one that eliminates additional memory overhead while enabling efficient GEMM execution in both FP16 and FP8 modes.
+An alternative approach is to store only the FP16 model and perform on-the-fly quantization to perform computation in FP8.
+Unlike the separate storage strategy, this approach avoids additional memory capacity overhead.
+However, its FP8 performance is suboptimal; although computation is carried out by FP8 units with high throughput, the full 16-bit weights must still be loaded from memory, effectively doubling memory traffic compared to native FP8 execution.
+As a result, memory bandwidth becomes a bottleneck, limiting the achievable throughput and diminishing much of FP8’s advantage over the FP16 baseline.
+These limitations highlight the need for a more sophisticated dual-precision serving mechanism—one that eliminates additional memory overhead while enabling efficient GEMM execution in both FP8 and FP16 modes.
 4 NestedFP
-4.1 NestedFP Overview
-For efficient dual-precision LLM serving, we propose NestedFP, a scheme that stores only 16-bit weights in memory and extracts 8-bit upper tensor from them when operating in FP8 mode.
-This approach not only avoids additional memory usage, but also prevents memory traffic amplification—such as fetching 16-bit weights when performing 8-bit GEMM.
-Figure 2 illustrates the core concept of NestedFP, which addresses the following two key challenges to make this approach viable.
-First, the extracted 8-bit representation must maintain accuracy comparable to existing FP8 quantization methods.
-Naively truncating the upper 8 bits of FP16 values (which use the E5M10 format) produces a low-precision representation similar to E5M2, which offers limited precision compared to the commonly preferred E4M3 format for LLM inference.
-Note that FP16 is basically E5M10.
-Moreover, this naive truncation often performs worse due to indiscriminate rounding, further degrading accuracy.
-This motivates the need for a quantization-aware format that simultaneously supports both FP16 and FP8 execution.
+Overview.
+Figure 2 illustrates the core concept of NestedFP, which stores only 16-bit weights in memory and extracts 8-bit weights from them when operating in FP8 mode.
+Specifically, NestedFP decomposes each FP16 parameter into two parts: the upper 8 bits and the lower 8 bits.
+With these two parts of the weight parameters stored as separate tensors, both parts are loaded in FP16 mode, whereas only the upper part is loaded in FP8 mode.
+This approach not only eliminates additional memory usage but also prevents memory traffic amplification—for example, fetching 16-bit weights when executing in 8-bit mode.
+Challenge 1: 8-bit Mode Accuracy.
+One key challenge of this approach is ensuring that the 8-bit representation extracted from 16-bit parameters achieves sufficient accuracy.
+In other words, it must maintain performance comparable to existing FP8 quantization methods.
+Naive truncation (simply using the upper 8 bits of FP16 values) yields a representation similar to E5M2, which offers limited precision compared to the commonly preferred E4M3 format for LLM inference.
+Moreover, such truncation often performs worse due to indiscriminate rounding, further degrading accuracy.
+This motivates the need for a quantization-aware data format that enables the extraction of high-quality FP8 representations.
 Section 4.1 elaborates on this data format design.
-Second, kernel efficiency must be maintained across both precisions.
-Achieving this requires storing the two 8-bit halves of each 16-bit weight as separate tensors.
-On H100 GPUs, where DRAM is accessed in 64-byte units, extracting individual bytes from packed 16-bit data results in wasted bandwidth.
-Separating the 8-bit components enables efficient FP8 execution using a single tensor, while FP16 models require reconstructing the original 16-bit weights via a custom GEMM kernel that dynamically fuses the components at runtime.
+Challenge 2: 16-bit Mode Throughput.
+Another key challenge is maintaining throughput in the 16-bit mode.
+The 8-bit mode in NestedFP trivially achieves high performance, as native FP8 GEMM kernels can be used directly after loading the upper-part tensor.
+However, the 16-bit mode is less straightforward: both the upper and lower-part tensors must be loaded and combined to reconstruct the original 16-bit weights.
+This requires a custom GEMM kernel capable of performing the reconstruction dynamically.
 Section 4.2 details this kernel design.
-Figure 2: FP16 weight is first pre-processed into 8-bit Upper and Lower parts. Both parts are used for FP16 model execution while only Upper part is used for FP8 model execution.
+Figure 2: FP16 weight is decomposed into 8-bit upper and lower parts. Both parts are used for FP16 mode while only upper part is used for FP8 mode.
 4.1 Compact yet Effective Data Format for Dual-Precision LLM
 Opportunities: Low-Entropy Exponent Bit in FP16 Models.
-Figure 3 shows the the weight distributions of FP16 linear layers—including QKVO projections and MLPs—across four different LLMs.
+Figure 3 shows the the weight distributions of all FP16 linear layers across four different LLMs.
 The vast majority of weight values have absolute magnitudes less than or equal to 1.75, which means that their most significant exponent bits in FP16 are zero.
@@
 This strong tendency suggests an opportunity for lossless exponent mapping from FP16 to the FP8 E4M3 format, which also uses 4 exponent bits.
-Min Max NestedFP Applicable Layers
-Llama 3.1 (8B) -0.92 0.91 128 of 128 (100%)
-Mistral Nemo Base (12B) -0.92 0.95 160 of 160 (100%)
-Phi-4 (14B) -3.60 3.89 146 of 160 (91.25%)
-Mistral Small Base (24B) -0.49 0.45 160 of 160 (100%)
-Figure 3: (a) Weight distributions of linear layers across different models. (b) Per-model weight range (min/max) and count of linear layers eligible for NestedFP quantization.
+Min Max NestedFP Applicable Layers Llama 3.1 (8B) -0.92 0.91 128 of 128 (100%) Mistral Nemo Base (12B) -0.92 0.95 160 of 160 (100%) Phi-4 (14B) -3.60 3.89 146 of 160 (91.25%) Mistral Small Base (24B) -0.49 0.45 160 of 160 (100%)
+Figure 3: (a) Weight distributions of linear layers across different models. (b) Per-model weight range (min/max) and the proportion of linear layers where NestedFP can be applied.
 Our Proposal.
-Leveraging the unused exponent bit in FP16, NestedFP restructures FP16 values so that high-quality E4M3 representations can be directly extracted.
-Figure 4a illustrates the refactoring process.
-The lower 8 bits of new format of NestedFP is simply constructed by inheriting lower 8 bits of the FP16 mantissa.
-The upper 8 bits are formed from the sign bit and 4 exponent bits (excluding the most significant bit of the FP16 exponent) of the original FP16 value, along with a 3-bit mantissa obtained by quantizing its 10-bit mantissa.
-This simple mapping scheme introduces a fixed scaling factor determined by the difference in exponent bias between the FP16 and E4M3 formats, which is $2^{8}$.
-In other words, NestedFP effectively applies E4M3 quantization with a scaling factor of $2^{8}$.
-To quantize the 10-bit FP16 mantissa down to 3 bits, we employ a round-to-nearest-even strategy.
-Specifically, we examine the 7 least significant bits of the FP16 mantissa (i.e., bits M[4:10]) to determine whether the value lies above or below the midpoint (64).
-If the value is greater than 64, we round up.
-If the value equals 64, we round up only when the least significant bit of the resulting 3-bit mantissa (M3) is 1.
-If the value is less than 64, we round down.
-This division process is done offline prior to model execution.
-FP8 GEMM can be performed directly using the upper 8 bits, whereas FP16 GEMM requires online reconstruction of 16-bit values, introducing additional complexity.
-Figure 4b illustrates this reconstruction process.
-Essentially, it concatenates the upper 8 bits and lower 8 bits, but not by simple concatenation.
-The earlier rounding decision must be reversed by checking an implicit checksum; if the LSB of the upper 8 bits differs from the MSB of the lower 8 bits, a rounding-down correction is applied before recombining the exponent and mantissa bits into a valid FP16 value.
-This process ensures precise, lossless recovery of the original FP16 values.
+Figure 4 illustrates how NestedFP decomposes an FP16 weight such that the upper 8 bits can serve as a high-quality E4M3 representation by leveraging an unused exponent bit in FP16.
+First, the lower 8 bits directly inherit the lower 8 bits of the FP16 mantissa.
+The upper 8 bits, which correspond to the E4M3 representation, are constructed as follows.
+The sign bit is directly inherited from FP16.
+The next four exponent bits are also copied from FP16, excluding the most significant exponent bit, which is assumed to be unused (or zero).
+Finally, 3 most significant mantissa (i.e., bits M[1:3]) bits are obtained and constructed value is applied rounding using a round-to-nearest-even policy.
+Specifically, the 7 least significant bits of the FP16 mantissa (i.e., bits M[4:10]) are examined to determine whether the value lies above or below the midpoint (64).
+If the value is greater than 64, it is rounded up; if it equals 64, it is rounded up only when the least significant bit of the resulting 3-bit mantissa (M3) is 1; otherwise, it is rounded down.
+This decomposition process is performed offline prior to model execution.
+In this scheme, removing most significant exponent bit effectively acts as multiplying $2^{8}$ between E4M3 and FP16 value.
+In other words, NestedFP effectively applies E4M3 quantization with a fixed scaling factor of $2^{8}$.
+FP16 Reconstruction.
+Figure 4 illustrates how NestedFP reconstructs FP16 weights at runtime.
+Essentially, this process merges the upper and lower 8-bit parts, but not through a simple concatenation.
+The rounding decision made during decomposition must be reversed by checking an implicit checksum; if the LSB of the upper 8 bits differs from the MSB of the lower 8 bits, a rounding-down correction is applied before recombining the exponent and mantissa bits into a valid FP16 value.
+This process ensures precise, lossless recovery of the original FP16 weights.
 (a) Refactoring to NestedFP
 (b) Reconstruction of FP16
-Figure 4: (a) Pre-processing FP16 weight value into two 8-bit NestedFP components (Offline). (b) On-the-fly reconstruction of FP16 weight values for tensor core execution (Online).
+Figure 4: (a) Decomposition of FP16 weight values into two 8-bit parts (offline). (b) Reconstruction of FP16 weight values (online).
 Handling Exception Layers.
-Since most models do not contain weights with absolute values exceeding 1.75, we leave these values in their original FP16 format without pre-processing.
-As a result, the corresponding layers are always executed in FP16, consistent with standard FP16 model execution.
-As shown in Sections 5, this exception handling preserves both model accuracy and end-to-end performance, as these layers constitute only a small portion of the total linear layers.
-4.3 GPU Kernel Design
-While executing FP8 GEMM in NestedFP is straightforward—simply performing GEMM on the NestedFP upper tensor, FP16 GEMM is more complex.
-Figure 6 illustrates the FP16 GEMM process in both a standard FP16 CUTLASS kernel and the NestedFP16 kernel.
-In the standard FP16 kernel, weights are loaded into shared memory, transferred to registers, and directly consumed by tensor cores without any transformation.
-In contrast, the FP16 path in NestedFP requires a reconstruction step before tensor core execution; weights are stored as two 8-bit tensors, which must be loaded into shared memory, copied to registers, and then reconstructed into full FP16 values through SIMT-level bitwise operations.
+For layers containing weight values with magnitudes exceeding 1.75 (i.e., where the MSB of the exponent is non-zero), we retain them in FP16 and always perform FP16 computation.
+This may reduce the benefits of dual-precision LLM serving, as the throughput advantage of FP8 mode becomes less pronounced.
+However, it is important to recall that, as shown in our analysis of popular models (Figure 3), such layers are rare and therefore have minimal impact on the overall effectiveness of the dual-precision scheme.
+4.2 FP16 GEMM Kernel with On-the-fly Reconstruction
+Figure 5 illustrates the high-level execution flow of the NestedFP FP16 GEMM kernel compared to a standard FP16 GEMM kernel.
+We assume an NVIDIA GPU environment, where the CUTLASS implementation serves as the standard FP16 GEMM kernel, upon which our kernel is also built.
+In the standard FP16 GEMM kernel, weights are loaded into shared memory, transferred to registers, and directly consumed by compute units (e.g., tensor cores in NVIDIA GPUs) without any transformation.
+In contrast, the FP16 GEMM in NestedFP requires a reconstruction step before tensor core execution; weights are stored as two 8-bit tensors, which must be loaded into shared memory, copied to registers, and then reconstructed into full FP16 values.
 Only after this reconstruction can the data be used by the tensor cores.
-This additional step introduces a source of overhead absent in traditional FP16 pipelines.
-To prevent it from becoming a performance bottleneck, SIMT operations must be highly efficient, and their latency must be effectively hidden within the broader compute pipeline.
-We present two key optimization strategies to address this challenge: (1) SIMT operation optimization and (2) pipelining and scheduling optimizations to prevent SIMT operations from stalling the tensor core pipeline.
-SIMT Reconstruction Operation Optimization.
-To mitigate the overhead introduced by SIMT operations, we fuse four 8-bit bitwise operations into a single 32-bit operation, following a strategy similar to any_precision_llm ; fp6_llm .
-As shown in Figure 6, the reconstruction process combines two 8-bit tensors, $W_{1}$ and $W_{2}$, stored in registers, into a full 16-bit FP16 weight.
-It begins with a correction step for $W_{1}$, followed by dropped $E_{1}$ restoration, and ends with the final fusion of $W_{1}$ and $W_{2}$.
-In this scheme, the lower 8 bits of the FP16 mantissa, $M[3:10]$, are taken directly from $W_{2}$.
-The upper tensor, $W_{1}$, contains $E[2:5]$ and $M[1:2]$, which may require correction depending on rounding.
-Rounding is detected using a checksum condition: the LSB of $W_{1}$ ($M_{3}^{\prime}$) and the MSB of $W_{2}$ ($M_{3}$) should match.
-If $M_{3}^{\prime}=1$ and $M_{3}=0$, no correction is needed for $E[2:5]$ and $M[1:2]$.
-If $M_{3}^{\prime}=0$ and $M_{3}=1$, rounding may have corrupted $E[2:5]$ and $M[1:2]$, and correction is applied by subtracting 1 at the $M_{3}$ bit position.
-To unify both cases and eliminate branching, we apply a branch-free correction; by subtracting $M_{3}$ from $W_{1}$ and taking only bits that corresponds to bit position $E[2:5]$ and $M[1:2]$ from $W_{1}$.
-Omitted zero is then restored and the full FP16 value is reconstructed using NVIDIA’s __byte_perm instruction.
+Not to make this additional step of reconstruction introduce a bottleneck, we present two key optimizations: (1) merged bitwise operations and (2) a three-stage pipeline.
+Efficient FP16 Reconstruction via Merged Bitwise Operations.
+FP16 reconstruction involves a large number of 8-bit bitwise operations.
+To reduce the associated overhead, we fuse four 8-bit operations into a single 32-bit instruction, following the approach in any_precision_llm ; fp6_llm .
+Algorithm 1 illustrates how FP16 reconstruction is performed using merged bitwise operations.
+Conceptually, the process iterates over the weight tensor, where each iteration processes four elements.
+In each iteration, four 8-bit values are loaded from the upper and lower tensors ($W_{\text{upper}}$ and $W_{\text{lower}}$) into 32-bit registers u and l, respectively (Line 5–6).
+Then, the sign bits are extracted from u and stored in another register s (Line 7).
+Next, the algorithm checks whether rounding has occurred by examining l (Line 8); if rounding is detected, the effect is reversed, thereby recovering the exponent and mantissa parts of the upper 8 bits (Line 9).
+Then, the original upper 8 bits, u_orig, are fully reconstructed by concatenating them with s (Line 10).
+Finally, u_orig and l are fused using NVIDIA’s __byte_perm instruction (lines 11–12), completing the reconstruction.
+Note that all the above operations are simultaneously applied to four 8-bit values through a 32-bit instruction, effectively reducing the total number of bitwise operations required.
 Figure 5: FP16 execution flow comparison.
-Figure 6: Algorithm of SIMT-Efficient Bitwise Reconstruction of FP16 Weights.
-Pipelining and Scheduling Optimizations.
-The standard FP16 GEMM kernel sources weight elements from registers and employs a two-stage pipeline, overlapping shared memory-to-register transfers with tensor core execution.
-Following prior work that integrates SIMT operations into the tensor core pipeline fp6_llm , we extend this design to a three-stage pipeline that concurrently overlaps register to shared memory transfers, SIMT operations, and tensor core execution.
-However, simply adopting a three stage pipeline is insufficient to fully hide the overhead of SIMT operations.
-Through empirical analysis, we find that distinct optimization strategies are required for non-cooperative and cooperative kernels.
-In the non-cooperative kernel, four warps jointly execute a single warp group MMA (Matrix Multiply Accumulate), whereas the cooperative kernel employs eight warps to execute two warp group MMAs concurrently.
-While shared memory transfers and tensor core executions are asynchronous operations—allowing subsequent instructions to be issued without waiting for completion—SIMT operations are synchronous and block instruction issue during their execution.
-In the non-cooperative kernel, we advance the scheduling of shared memory to register transfers relative to the naive three stage pipeline and issue multiple copies in bulk.
-This approach not only allows multiple shared memory to register copy concurrent with SIMT execution, but also allows SIMT instruction to be issued without delay as operands are pre-loaded into registers.
-However, in the cooperative kernel, which leverages additional compute parallelism by executing two warp group MMAs concurrently, warp groups contend for shared resources such as SIMT pipelines, tensor core pipelines, and shared memory bandwidth.
-This contention reduces the effectiveness of the optimizations used in the non-cooperative kernel.
-To address this, we explored several alternatives and found that inserting a compiler-level fence—specifically, an NVVM code motion barrier—after the transformation stage significantly improves performance.
-This fence inhibits compiler optimizations that may delay or reorder memory operations and serves as a signal for the imminent use of a transformed input in a register, preventing the compiler from eliminating or deferring its materialization.
+Algorithm 1 FP16 Weight Reconstruction
+[…]
+2WW: reconstructed FP16 weight tensor
+3NN: total number of elements in the weight tensor
+4for $i=0$ to NN step 44 do
+5uint32_t u $W_{\text{upper}}[i:i+4]$
+6uint32_t l $W_{\text{lower}}[i:i+4]$
+7uint32_t s $u\mathbin{\&}\texttt{0x80808080}$
+8uint32_t sub $(l\mathbin{\&}\texttt{0x80808080})\gg 7$
+9uint32_t u_orig $(((u-sub)\gg 1)\mathbin{\&}\texttt{0x3f3f3f3f})$
+10u_orig $\texttt{u\_orig}\mathbin{|}\texttt{s}$
+11W[i:i+2]W[i:i+2] $byte\_perm(\texttt{u\_orig},l,\texttt{0x1504})$
+12W[i+2:i+4]W[i+2:i+4] $byte\_perm(\texttt{u\_orig},l,\texttt{0x3726})$
+13end for
+Three-Stage Pipeline.
+The CUTLASS FP16 GEMM kernel, upon which the NestedFP FP16 kernel is built, employs a two-stage pipeline to overlap data transfer and computation.
+Specifically, the two stages correspond to (1) data movement between shared memory and registers and (2) Tensor Core execution, which are overlapped for high throughput.
+NestedFP extends this design into a three-stage pipeline to further hide reconstruction overheads.
+In particular, the three stages in NestedFP are: (1) shared-memory-to-register data transfer, (2) FP16 reconstruction, and (3) Tensor Core execution.
 5 Evaluation
-5.1 NestedFP8 Accuracy
-To evaluate the accuracy of our NestedFP8 model, we compare it with an FP8 model using commonly adopted quantization configurations—per-channel weight quantization and per-token activation quantization, both scaled by the absolute maximum value.
-We evaluate on four models (Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), and Mistral Small(24B)) using three downstream tasks: Minerva Math, MMLU Pro, and BBH.
-All models are integrated into the vLLM framework, and accuracy is assessed using the LM Evaluation Harness lm-evaluation-harness framework.
-As noted in Section 4.2, 8.75% of linear layers in Phi-4 are not applicable to NestedFP and are thus executed in FP16.
-Our method internally applies a single global scaling factor (256) to all weight tensors.
-For activation quantization, we apply per-tensor scaling based on the absolute maximum value.
-As shown in Table 2, our method achieves accuracy comparable to the FP8 baseline despite foregoing fine-grained quantization, demonstrating that NestedFP8 can preserve accuracy.
-Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.9 16.8 42.9 41.3 34.6 35.7 MMLU Pro 32.8 33.0 34.9 35.0 52.7 53.3 53.6 53.7 BBH 38.6 38.8 41.2 40.6 27.1 27.9 51.8 51.4
-Table 2: Accuracy comparison between FP8 baseline and NestedFP8. FP8(B) denotes FP8 baseline, and FP8(N) denotes NestedFP8.
-5.2 NestedFP16 Kernel Performance
-Comparison over Vanilla FP16 Kernel.
-To compare the performance of NestedFP16 kernel with normal FP16 GEMM kernel fairly, we implemented a highly optimized FP16 baseline using the same CUTLASS codebase.
-This baseline selects the best configuration for each GEMM shape through exhaustive search over the design space, closely matching cuBLAS performance (validated in Appendix).
-Our kernel design space includes non-cooperative and cooperative variants.
-For non-cooperative kernels, we perform a grid search over tile shapes with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$.
-Cooperative kernels use larger tiles ($T_{n}\in{128,256}$), and configurations that fail to compile are excluded.
-Non-cooperative kernels use $(1,1,1)$ clusters, while cooperative ones use $(2,1,1)$, assigning two thread blocks per GPU Processing Cluster.
-We use the standard data parallel scheduler for non-cooperative kernels and both the standard data parallel and Stream-K streamk schedulers for cooperative ones.
-All benchmarks are conducted on an NVIDIA H100 SXM GPU.
-Our NestedFP16 kernels are optimized using the same design space as the FP16 baseline, ensuring that each GEMM shape is paired with its optimal kernel configuration.
-We benchmark both kernels across all linear-layer GEMM shapes used in LLaMA 3.1–8B, Mistral Nemo, Phi-4, and Mistral Small, each of which involves four distinct $(N,K)$ shapes.
-The batch dimension $(M)$, determined by the number of input tokens during inference, is swept from 32 to 2048 in increments of 32, yielding a total of 882 GEMM configurations across 14 unique $(N,K)$ shapes.
-Figure 7 compares the performance of the largest $(N,K)$ GEMM configurations across the four models.
-As shown, our kernel consistently maintains low SIMT overheads across all $M$ values.
-Compared to the CUTLASS baseline, the average relative overheads across all GEMM configurations of each model are 5.69% (LLaMA 3.1–8B), 6.83% (Mistral Nemo), 6.65% (Phi-4), and 6.71% (Mistral Small).
-Additional results for other GEMM shapes, as well as comparisons with PyTorch W16A16 kernels, which exhibit similar trends, are provided in Appendix.
-While certain kernel performance regarding certain $M$ show more variability, the average overhead remains modest, demonstrating the practicality of our approach across diverse shapes in real LLM inference scenario.
-Impact of Optimization Techniques.
-To quantify the impact of successive kernel-level optimizations, we measure latency for a matrix multiplication of shape $M\times 5120\times 32768$.
-Figure 7 illustrates the latency reduction achieved across three optimization stages.
-Level 1 introduces a three-stage pipeline that overlaps shared memory transfers with SIMT execution and tensor core operations.
-Level 2 incorporates SIMT-level FP16 reconstruction optimization.
-This transition yields a substantial 38.3% reduction in latency relative to Level 1.
-Level 3 applies additional pipelining and scheduling optimizations resulting in a further 11.0% reduction compared to Level 2.
-Each stage contributes a measurable improvement, highlighting the cumulative benefits of the proposed optimization strategies.
-Figure 7: (a) CUTLASS baseline and NestedFP16 kernel performance comparison. (b) Kernel latency reduction with different levels of optimization strategies.
-5.3 End-to-End Performance
-We integrated NestedFP into the vLLM framework as a quantization configuration.
-The system supports per-iteration precision switching, allowing the scheduler to dynamically select FP16 or FP8 based on workload and SLO demands.
-To ensure fair comparison, both FP16 baselines and NestedFP models were compiled using the vLLM 0.8.3 V1 engine and PyTorch 2.6.0+cu12.4.
-Experiments were run on a single NVIDIA H100 GPU with fixed request sizes (256 input, 512 output tokens) and batch sizes ($M$) ranging from 32 to 512 to assess throughput scaling.
-Throughput is reported in tokens per second (tok/s).
-Figure 8 shows end-to-end performance under FP16 and NestedFP.
-The FP16 runs use torch.matmul, which dispatches to cuBLAS by default within the vLLM framework.
-Compared to this baseline, NestedFP16 introduces only minor overhead—4.50% (LLaMA 3.1–8B), 4.51% (Mistral Nemo), 2.69% (Phi-4), and 3.91% (Mistral Small)—generally lower than kernel-level overheads, likely due to amortization across non-GEMM components.
-In contrast, NestedFP8 achieves clear speedups over NestedFP16: 1.24$\times$, 1.32$\times$, 1.29$\times$, and 1.53$\times$, respectively.
-Larger models gain more, reflecting the growing cost of matrix multiplication.
-Overall, NestedFP improves efficiency for large-scale LLMs.
-Figure 8: End-to-end performance comparison with four models using FP16, NestedFP16 and NestedFP8.
+Through extensive experiments, we demonstrate that NestedFP is an effective solution for handling fluctuating loads in LLM serving by dynamically switching between FP16 and FP8 modes, as evidenced by the following three key results:
+  • The quality of FP8 model extracted from FP16 parameters in the NestedFP data format matches state-of-the-art FP8 quantization results. (Section 5.1)
+  • The FP16 mode of NestedFP achieves throughput comparable to baseline FP16 inference using the vanilla CUTLASS kernel, despite performing on-the-fly FP16 reconstruction.
+(Section 5.2 and Section 5.3)
+  • NestedFP significantly improves SLO attainment over FP16-only deployment, without introducing additional memory overhead. (Section 5.4)
+5.1 FP8 Model Quality
+Methodology.
+We evaluate the accuracy of the FP8 model in NestedFP against FP8 models quantized using commonly adopted configurations, specifically per-channel weight quantization.
+For activation quantization, we apply per-token quantization for both NestedFP and the baseline.
+The evaluation covers six models: Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), Mistral Small (24B), Llama 3.1 70B, and DeepSeek-R1-Distill-Llama-70B deepseek_ref ; nemo_ref ; small_ref ; llama3_technical_report ; phi_ref .
+For benchmarks, we use three downstream tasks: Minerva Math, MMLU Pro, and BBH minerva ; bbh ; mmlupro .
+All models are integrated into the vLLM framework, and accuracy is measured using the LM Evaluation Harness lm-evaluation-harness .
+As noted in Section 4.1, 8.75% of the linear layers in Phi-4 are not applicable to NestedFP and are therefore executed in FP16.
+More details on the experimental setup are provided in Appendix A.
+Results.
+Table 2 presents the results.
+Across all models and benchmarks, the FP8 model in NestedFP achieves accuracy comparable to the baseline FP8 model, demonstrating the effectiveness of the new data format of NestedFP.
+Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small Llama 3.1 70B DeepSeek-R1- Distill-Llama-70B FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.5 16.4 42.9 42.9 34.2 35.2 35.7 34.6 46.0 45.5 MMLU Pro 32.8 33.2 35.3 34.7 52.7 53.3 53.6 53.7 48.1 47.4 49.2 49.4 BBH 38.6 39.0 40.6 40.0 27.1 28.7 50.8 51.1 49.1 49.7 53.2 53.4
+Table 2: Accuracy of FP8 models in NestedFP across various downstream tasks, compared to the baseline FP8 model. FP8(B) denotes the baseline model, and FP8(N) denotes NestedFP.
+5.2 FP16 GEMM Kernel Performance
+Methodology.
+We evaluate the FP16 GEMM kernel performance of NestedFP against the CUTLASS FP16 GEMM kernel, upon which ours is built, using an NVIDIA H100 GPU.
+Note that NestedFP does not require any custom kernel for FP8 GEMM; thus, we omit FP8 kernel evaluation.
+We consider four different weight matrix shapes: (N, K) = (28672, 4096), (28672, 5120), (35840, 5120), (65536, 5120).
+The $(M)$ dimension, which corresponds to the number of tokens processed in parallel, is varied from 32 to 2048 in increments of 32.
+For both kernels, we select the best configuration (e.g., tile size) for each GEMM shape through an exhaustive search over the design space.
+Details of the kernel search space are provided in Appendix D.2.
+Results.
+Figure 6 shows the results.
+The NestedFP kernel consistently demonstrates performance comparable to the CUTLASS baseline across all GEMM shapes.
+Specifically, the average relative overheads for each weight matrix dimension are 5.89%, 6.33%, 4.59%, and 6.33%, respectively.
+These results indicate that the NestedFP kernel effectively minimizes the overhead introduced by on-the-fly FP16 reconstruction.
+Additional results for other GEMM shapes, which exhibit similar trends, as well as comparisons with cuBLAS FP16 kernels, are provided in Appendix B and Appendix D.1.
+Figure 6: CUTLASS baseline and NestedFP kernel performance comparison.
+5.3 FP16 Mode End-to-End Inference Performance
+Methodology.
+We evaluate the end-to-end inference throughput of the FP16 mode of NestedFP by integrating it into the vLLM framework on a single NVIDIA H100 GPU.
+Specifically, we set the input and output sequence lengths to 1024 and 512, respectively, while varying the batch size from 32 to 512.
+More details on the experimental setup are provided in Appendix C.
+The results are compared against vanilla vLLM FP16 execution, which relies on PyTorch for invoking GEMM kernels.
+Again, the FP8-mode throughput of NestedFP is identical to baselines such as vanilla vLLM and is therefore omitted from evaluation.
+Results.
+Figure 7 presents the results.
+NestedFP achieves FP16-mode throughput highly comparable to the baseline, with only minor overheads—5.04% (LLaMA 3.1–8B), 5.56% (Mistral Nemo), 4.76% (Phi-4), and 4.56% (Mistral Small) on average across batch sizes.
+The end-to-end throughput degradation is even lower than the kernel-level overheads reported in Section 5.2, since non-GEMM components also contribute to total execution time.
+Additional end-to-end throughput results are provided in Appendix C.
+Figure 7: FP16 inference throughput comparison of four models between vanilla vLLM and NestedFP.
+5.4 Impact on SLO Attainment
+Methodology.
+In this section, we present a case study demonstrating how NestedFP can improve SLO attainment.
+We implement a simple load-aware mode switching policy for NestedFP that, in each iteration, selects FP8 mode when the number of tokens to process exceeds a predefined threshold (1024 in this experiment); otherwise, it selects FP16 mode.
+The evaluation is conducted using Llama 3.1–70B on a cluster equipped with four NVIDIA H100 GPUs.
+As input, we use 1,000 requests sampled from an Azure LLM inference trace while varying the request rate.
+We measure p90 TTFT, p90 TPOT, and SLO attainment under two criteria: loose and tight.
+The loose SLO is defined as $\leq 10\times$ the average single-request latency, and the tight SLO is defined as $\leq 5\times$ the average single-request latency, following the methodology of a previous study sola_llm .
+Results.
+Table 3 shows the results.
+NestedFP consistently outperforms the FP16-only baseline across all request rates in terms of TTFT, TPOT, and SLO attainment, with particularly pronounced improvements under high request rates.
+Note that this experiment employs a simple, straightforward switching policy, leaving substantial room for further optimization by integrating NestedFP with more sophisticated policies—such as those considering KV-cache utilization or incorporating load prediction mechanisms.
+Nevertheless, these results confirm that NestedFP provides an efficient dual-precision mechanism that can significantly enhance SLO attainment in LLM serving systems.
+Request Rate Method Tight SLO (%) Loose SLO (%) p90 TTFT (s) p90 TPOT (ms)
+7.47 req/s FP16 0.7 14.1 9.9893 162.2
+NestedFP 72.4 90.0 1.9518 118.9
+5.03 req/s FP16 71.6 93.6 1.6449 151.3
+NestedFP 94.6 99.9 0.9679 86.8
+3.79 req/s FP16 83.1 96.9 1.2404 109.2
+NestedFP 96.6 100 0.8168 60.1
+1.99 req/s FP16 90.4 99.4 0.8915 64.9
+NestedFP 98.4 100 0.6116 47.7
+Table 3: SLO attainment and p90 TTFT/TPOT of Llama-3.1-70B on 4×H100 under varying loads.
 6 Conclusion
-In this paper, we introduce NestedFP, a precision-aware serving framework designed to better satisfy service-level objectives (SLOs) under dynamic request conditions.
-We propose a novel approach for decomposing FP16 weights into two 8-bit tensors, enabling one to be used directly as an FP8 representation.
-Our FP8 format achieves accuracy comparable to that of standard quantized FP8 models.
-In addition, we develop a custom 16-bit GEMM kernel tailored to this dual-format representation.
-Experimental results show that our kernel achieves an average performance gap of 6.47% compared to the CUTLASS baseline kernel.
-Furthermore, by integrating our method into the vLLM framework, we observe up to 1.55$\times$ speedup in FP8 model execution relative to FP16 model execution, while incurring average 3.90% overhead across models when executing in FP16 mode.
-These results demonstrate that NestedFP successfully bridges the gap between FP16 accuracy and FP8 efficiency, requiring the same memory footprint as a standard FP16 model while supporting both precision modes.
-By enabling iteration-level precision selection, our approach allows production LLM serving systems to dynamically balance performance and accuracy requirements under varying workload conditions, providing a flexible and efficient solution for real-world deployment scenarios.
+We introduce NestedFP, a memory-efficient framework that supports both FP16 and FP8 precisions for LLM inference in a memory-efficient way.
+NestedFP introduces a new data format that enables direct extraction of FP8 weights from FP16 weights without additional memory overhead.
+NestedFP also incorporates a custom GEMM kernel optimized for this format to ensure efficient computation.
+As a result, NestedFP enables dynamic switching between FP16 and FP8 modes with the memory footprint of FP16, without noticeable degradation in either accuracy or inference throughput, providing an effective means to improve SLO attainment in LLM serving.
+7 Acknowledgments
+This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea Government (MSIT) (RS-2024-00340008) and Institute of Information & Communications Technology Planning & Evaluation (IITP) under the artificial intelligence semiconductor support program (IITP-2023-RS-2023-00256081), funded by the Korea Government (MSIT).
 Appendix A Evaluation Details
@@
 Kernel performance was measured using NVIDIA Nsight Compute.
-In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$ as it provides more robust performance.
+In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$.
 During performance measurement, we flushed all caches using the cache-control option, but did not fix the GPU clock frequency using the clock-control option.
-Appendix B Extended Kernel Performance for NestedFP16
-Figure 9: Performance comparison between our CUTLASS baseline and NestedFP16 kernel across 14 unique $(N,K)$ GEMM shapes.
+Appendix B Extended Kernel Performance for NestedFP
+Figure 8: Performance comparison between our CUTLASS baseline and NestedFP kernel across 14 unique $(N,K)$ GEMM shapes.
 Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
@@
 This consistency highlights the effectiveness of our kernel-level optimizations.
-Appendix C Extended End-to-End Throughput Evaluation
+Appendix C Extended End-to-End Throughput Evaluation for NestedFP
 (a) (32, 512)
@@
 Figure 9: Extended end-to-end throughput evaluation across different input/output token configurations.
-Each subplot shows throughput comparison for FP16, NestedFP16, FP8 and NestedFP8 across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
+Each subplot shows throughput comparison for FP16 and NestedFP across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
 We extend our end-to-end evaluation to further demonstrate the robustness of NestedFP.
 Figure 9 presents results for four additional input/output configurations: (32, 512), (1024, 512), (32, 32), and (1024, 32).
-Across all settings, NestedFP16 exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 3.88% and a maximum of 8.22%.
-On a per-model basis, the average overhead is 5.15% for LLaMA 3.1 (8B), 4.19% for Mistral-Nemo (12B), 3.35% for Mistral-Small (24B), and 2.81% for Phi-4 (14B).
-In contrast, NestedFP8 achieves significant speedups over NestedFP16, with average throughput gains of 1.26$\times$ for LLaMA 3.1, 1.33$\times$ for Mistral-Nemo, 1.53$\times$ for Mistral-Small, and 1.36$\times$ for Phi-4 across all input/output shapes.
-Notably, the largest model, Mistral-Small, benefits the most from the NestedFP8 quantization.
-These results indicate that the performance gains of NestedFP are consistent and robust across a wide range of input/output configurations.
-When compared to Torch FP8, NestedFP8 maintains competitive performance, achieving 96.84% of Torch FP8’s throughput for LLaMA 3.1, 97.76% for Mistral-Nemo, 98.79% for Mistral-Small, and 96.77% for Phi-4.
-Our limited search space shows potential gain that can be achieved with further FP8 kernel optimization.
-We note that we did adopt Stream-K scheduling for NestedFP8.
+Across all settings, NestedFP exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 5.52% and a maximum of 7.69%.
+On a per-model basis, the average overhead is 5.97% for LLaMA 3.1 (8B), 6.67% for Mistral-Nemo (12B), 5.51% for Phi-4 (14B), and 5.05% for Mistral-Small (24B).
 Our evaluation setup mirrors the configuration used in the main throughput experiments.
-All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.3 V1 engine with PyTorch 2.6.0+cu12.4 for both the baseline and our proposed NestedFP implementations.
+Throughput was computed as the total number of generated tokens divided by the end-to-end latency for each test.
+All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.5 V1 engine with PyTorch 2.7.0+cu12.6 for both the baseline and our proposed NestedFP implementations.
 Chunked prefill, enabled by default in the vLLM V1 engine, was used throughout the evaluation.
@@
 At each iteration, the GEMM operation shape $(M,N,K)$ is determined based on the number of tokens grouped in that step, with $M$ denoting the total token count.
-To systematically evaluate throughput under different batch sizes, we adjusted the max_num_batched_token and max_num_seqs parameters to constrain the batch size accordingly.
+We set the maximum number of batched tokens to 8192 for the throughput test.
 Appendix D CUTLASS FP16 GEMM Kernel
-We extend the open-source NVIDIA CUTLASS 3.6 library to support our custom format.
-We begin by describing the efficient design of GEMM kernels within CUTLASS for the NVIDIA H100 GPU and introduce key components that define our design space.
-In the following section, we benchmark our CUTLASS kernels against PyTorch to validate strong performance of CUTLASS baseline.
-D.1 GEMM Kernel Design
-Figure 11: (a) Data parallel scheduling.
-A single output tile is processed by a single thread block.
-(b) Stream-K scheduling.
-A single output tile is processed collaboratively across multiple thread blocks.
-(c) DRAM to shared memory copy handled by 4 warps in producer warp group and TMA engine.
-Figure 12: (a) Non-cooperative kernel.
-A single warp group is responsible for processing entire input tile.
-(b) Cooperative kernel.
-Two warp groups are each responsible for processing half of the first input tile and corresponding output tile.
-GEMM operates on a weight matrix of shape $N\times K$ and an activation matrix of shape $M\times K$.
-To parallelize this computation, CUTLASS partitions the input matrices into tiles of size $T_{n}\times T_{k}$ and $T_{m}\times T_{k}$, respectively.
-These tiles define the granularity at which work is distributed across thread blocks.
-As illustrated in Figure 11, each input tile is processed along the $K$ dimension to produce a corresponding output tile, which is accumulated in registers.
-CUTLASS employs a persistent tile scheduler, assigning each thread block to a fixed streaming multiprocessor (SM) for the duration of execution.
-In data-parallel scheduling, each SM processes the full $K$ dimension for a given output tile independently.
-In contrast, stream-K scheduling enables multiple thread blocks to collaboratively compute a single output tile by dividing the $K$-dimension workload across thread blocks, as shown in Figure 11.
-The partial results are later merged to form the final output.
-In all cases, the computed output tiles are written back to global memory (DRAM).
-To hide DRAM latency and enable compute–memory overlap, tiles are asynchronously loaded from global memory into shared memory.
-As shown in Figure 11, shared memory is partitioned to hold multiple tiles, allowing future iterations to be staged while the current tile is being processed.
-On NVIDIA H100 GPUs, this transfer is managed by the Tensor Memory Accelerator (TMA) engine.
-CUTLASS uses warp specialization to pipeline these operations: one warp group (four warps) handles TMA data transfers (producers), while one or two other warp groups (consumers) execute matrix-multiply-accumulate (MMA) operations, depending on the kernel configuration.
-Once data is available in shared memory, the kernel proceeds with MMA using one of two configurations.
-In the RS (register-shared) kernel, weights are loaded into registers and processed with activations from shared memory—this approach enables bit-level manipulations required by our custom format.
-In contrast, SS (shared-shared) kernels source both operands directly from shared memory.
-On Hopper, SM90 MMA instructions require the second operand to come from shared memory.
-MMA operations are executed at the warp group granularity, requiring all warps in a group to be ready before issuing the instruction.
-Consequently, SIMT-level manipulations on weights must complete across all four warps before any warp group MMA operation begins.
-CUTLASS supports both non-cooperative and cooperative kernel modes.
-Non-cooperative kernels use a single consumer warp group, while cooperative kernels employ two consumer groups that share activations but process distinct weight tiles.
-The behavioral differences between these modes are illustrated in Figure 12.
-D.2 CUTLASS FP16 Baseline and cuBLAS W16A16 Kernel Comparison
-Figure 13: Performance comparison between PyTorch torch.matmul and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes.
+D.1 CUTLASS FP16 Baseline and PyTorch Kernel Comparison
+Figure 10: Performance comparison between PyTorch kernel and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes.
 Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
@@
 These results confirm that our CUTLASS baseline is not only competitive but often exceeds production-level kernel performance, ensuring a reliable and fair comparison when evaluating NestedFP’s overhead.
+D.2 Kernel Search Space
+The kernel design space consists of two variants: non-cooperative and cooperative kernels.
+For non-cooperative kernels, we perform a grid search over tile dimensions with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$.
+Cooperative kernels adopt larger tiles with $T_{n}\in{128,256}$, and configurations that fail to compile are excluded.
+Non-cooperative kernels use a cluster shape of $(1,1,1)$, whereas cooperative kernels use $(2,1,1)$.
+We employ the persistent scheduler for non-cooperative kernels, and both the persistent and Stream-K [25] schedulers for cooperative kernels.
 Appendix E Broad Applicability of NestedFP

### v3

@@
 1 Introduction
-Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information gpt4_technical_report ; gpt3_technical_report ; llama3_technical_report .
+Large Language Models (LLMs) have emerged as foundational components in modern AI systems, enabling a wide range of applications such as virtual assistants, text and code generation, and multi-modal tasks that integrate textual and visual information Achiam et al.
+(2023); Brown et al.
+(2020); Meta (2024).
 Their broad applicability and strong performance have driven widespread adoption, leading to substantial user demand.
-For example, OpenAI’s production-scale models process roughly 100 billion words per day openai_billion_token and production query rates frequently surpass 200 requests per second chatgpt_fact ; scale_llm .
-To accommodate such large-scale applications, research efforts have focused not only on improving serving throughput orca ; paged_attention but also on meeting service-level objectives (SLOs) sarathi_serve ; llumnix ; sola_llm .
+For example, OpenAI’s production-scale models process roughly 100 billion words per day Griffin (2024) and production query rates frequently surpass 200 requests per second Lammertyn ; Yao et al.
+(2024).
+To accommodate such large-scale applications, research efforts have focused not only on improving serving throughput Kwon et al.
+(2023); Yu et al.
+(2022) but also on meeting service-level objectives (SLOs) Agrawal et al.
+(2024); Hong et al.
+(2025); Sun et al.
+(2024).
 One of the key challenges in achieving SLOs for LLM serving lies in handling dynamic load.
@@
 As a promising alternative, we explore an approach we call dual-precision LLM serving, which utilizes FP8 together with FP16, the de facto standard number format for weight parameters.
-While FP8 may incur a slight degradation in model quality, it offers up to 2× higher peak throughput compared to FP16 fp8_formats_for_deep_learning ; training_and_inference_llm_fp8 ; efficient_ptq_fp8 .
+While FP8 may incur a slight degradation in model quality, it offers up to 2× higher peak throughput compared to FP16 Micikevicius et al.
+(2022); Perez et al.
+(2023); Shen et al.
+(2024).
 By operating in FP16 mode under normal conditions and switching to FP8 mode only during sudden load surges, we can effectively manage dynamic workloads without resource overprovisioning.
@@
 NestedFP decomposes each 16-bit weight tensor into two 8-bit components, allowing FP8 inference without additional memory overhead or throughput degradation.
-Since FP16 mode requires reconstructing 16-bit values by merging the two 8-bit components in this approach, we develop a custom General Matrix Multiplication (GEMM) kernel, built on the CUTLASS library cutlass , that performs accurate on-the-fly FP16 reconstruction during kernel execution.
-According to our evaluation, NestedFP enables high-quality, dual-precision LLM serving efficiently: FP8 models generated from decomposed FP16 weights match the quality of existing FP8 quantization methods, and our FP16 GEMM kernel incurs only a 6.47% average performance gap relative to the CUTLASS baseline, which translates to 4.98% end-to-end overhead across models.
+Since FP16 mode requires reconstructing 16-bit values by merging the two 8-bit components in this approach, we develop a custom General Matrix Multiplication (GEMM) kernel, built on the CUTLASS library NVIDIA , that performs accurate on-the-fly FP16 reconstruction during kernel execution.
+According to our evaluation, NestedFP enables high-quality, dual-precision LLM serving efficiently: FP8 models generated from decomposed FP16 weights match the quality of existing FP8 quantization methods, and our FP16 GEMM kernel incurs only a 6.38% average performance gap relative to the CUTLASS baseline, which translates to 4.98% end-to-end overhead across models.
 Our contributions are summarized as follows:
@@
 Commonly used 16-bit formats include FP16 (E5M10) and BF16 (E8M7), while prevalent 8-bit formats include E4M3 and E5M2.
-These formats are natively supported by many modern hardware platforms and have become standard choices in low-precision machine learning computations investigation_fp8_across_accelerators .
+These formats are natively supported by many modern hardware platforms and have become standard choices in low-precision machine learning computations Kim et al.
+(2025).
 2.2 FP8 Quantization for LLMs
@@
 First, floating-point quantization is generally better suited for LLMs than integer quantization due to its wider dynamic range.
-Unlike in computer vision tasks—where integer quantization is prevalent—LLMs often exhibit activation outliers amplified by operations such as LayerNorm efficient_ptq_fp8 .
-These outliers can significantly degrade model performance if not properly managed investigation_fp8_across_accelerators .
-Although recent work has attempted to mitigate this issue in integer quantization through static outlier-aware techniques efficient_ptq_fp8 ; smoothquant , these methods still underperform compared to floating-point quantization, which inherently better accommodates long-tailed distributions and high dynamic ranges llm_fp4 ; fp8_formats_for_deep_learning ; int_or_fp .
-Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs investigation_fp8_across_accelerators .
+Unlike in computer vision tasks—where integer quantization is prevalent—LLMs often exhibit activation outliers amplified by operations such as LayerNorm Shen et al.
+(2024).
+These outliers can significantly degrade model performance if not properly managed Kim et al.
+(2025).
+Although recent work has attempted to mitigate this issue in integer quantization through static outlier-aware techniques Shen et al.
+(2024); Xiao et al.
+(2023), these methods still underperform compared to floating-point quantization, which inherently better accommodates long-tailed distributions and high dynamic ranges Liu et al.
+(2023); Micikevicius et al.
+(2022); Zhang et al.
+(2024).
+Second, FP8 support is rapidly being adopted by modern hardware accelerators, including NVIDIA Hopper GPUs and Intel Gaudi HPUs Kim et al.
+(2025).
 Leveraging this support, FP8 GEMM operations can achieve up to $2\times$ speedup compared to their FP16 counterparts.
@@
 E4M3 Format.
-Among the two widely adopted FP8 formats—E4M3 and E5M2—existing studies have consistently shown that E4M3 yields higher inference accuracy for Transformer-based language models fp8_formats_for_deep_learning ; efficient_ptq_fp8 ; hybrid_fp8_training_and_inference .
-However, due to its reduced dynamic range compared to FP16, the effectiveness of E4M3 critically depends on the application of appropriate scaling strategies the_power_of_exponent ; fp8_formats_for_deep_learning ; hybrid_fp8_training_and_inference .
+Among the two widely adopted FP8 formats—E4M3 and E5M2—existing studies have consistently shown that E4M3 yields higher inference accuracy for Transformer-based language models Micikevicius et al.
+(2022); Shen et al.
+(2024); Sun et al.
+(2019).
+However, due to its reduced dynamic range compared to FP16, the effectiveness of E4M3 critically depends on the application of appropriate scaling strategies Kuzmin et al.
+(2022); Micikevicius et al.
+(2022); Sun et al.
+(2019).
 Scaling is applied to align the dynamic range of tensor values with the numerical bounds of the E4M3 format.
-Scaling factors may be determined statically—based on heuristics such as the absolute maximum, percentile clipping, or mean squared error (MSE) minimization fp8_formats_for_deep_learning —or computed dynamically during runtime.
-Weight tensors are typically scaled statically on a per-channel basis llm_fp4 ; efficient_ptq_fp8 ; int_or_fp , most commonly using the absolute maximum value.
-Activation tensors, by contrast, may be scaled offline on a per-tensor basis llm_fp4 ; efficient_ptq_fp8 ; int_or_fp or dynamically on a per-token basis during inference.
-While the latter approach can improve accuracy, it incurs additional runtime cost due to the overhead of computing scaling factors on-the-fly investigation_fp8_across_accelerators ; training_and_inference_llm_fp8 .
+Scaling factors may be determined statically—based on heuristics such as the absolute maximum, percentile clipping, or mean squared error (MSE) minimization Micikevicius et al.
+(2022)—or computed dynamically during runtime.
+Weight tensors are typically scaled statically on a per-channel basis Liu et al.
+(2023); Shen et al.
+(2024); Zhang et al.
+(2024), most commonly using the absolute maximum value.
+Activation tensors, by contrast, may be scaled offline on a per-tensor basis Liu et al.
+(2023); Shen et al.
+(2024); Zhang et al.
+(2024) or dynamically on a per-token basis during inference.
+While the latter approach can improve accuracy, it incurs additional runtime cost due to the overhead of computing scaling factors on-the-fly Kim et al.
+(2025); Perez et al.
+(2023).
 FP8 Results.
-While FP8 inference offers significant throughput gains, it comes with the potential caveat of quality degradation due to the inevitable loss of information associated with reduced precision dfloat11 .
+While FP8 inference offers significant throughput gains, it comes with the potential caveat of quality degradation due to the inevitable loss of information associated with reduced precision Zhang et al.
+.
 Table 1 presents a comparison between FP8 and FP16 across various models and benchmarks.
-Specifically, for FP8 quantization, we use the E4M3 format with per-token activation and per-channel weight quantization for all linear layers, using the vLLM framework vllm .
+Specifically, for FP8 quantization, we use the E4M3 format with per-token activation and per-channel weight quantization for all linear layers, using the vLLM framework vllm-project .
 A consistent accuracy degradation is observed in FP8 models, which clearly shows the trade-off between throughput and model quality in FP8 quantization.
@@
 3.1 Challenge: Dynamic Load in LLM Serving
-LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests andes ; hygen .
+LLM serving is characterized by substantial variability in request rates, as well as input prompt and output sequence lengths across requests Liu et al.
+(2024); Sun et al.
+(2025).
 These variability is particularly problematic because the batch size scheduled at each iteration can fluctuate significantly, directly affecting the large deviations in Time To First Token (TTFT) and Time Per Output Token (TPOT).
-Such fluctuations occur in a highly dynamic and often unpredictable manner llumnix , making it difficult for service providers to consistently meet SLOs.
-To quantify the extent of this variability, we analyze production traces from Microsoft Azure’s LLM inference service azure_llm_inference_trace .
+Such fluctuations occur in a highly dynamic and often unpredictable manner Sun et al.
+(2024), making it difficult for service providers to consistently meet SLOs.
+To quantify the extent of this variability, we analyze production traces from Microsoft Azure’s LLM inference service Azure (2024).
 Specifically, we measure how the request rate fluctuates over a one-hour interval (00:00–01:00 UTC, May 10, 2024, in the trace).
@@
 Moreover, as shown in the one-minute zoom-in view, these fluctuations occur at high temporal frequency, indicating second-level variability.
-A classic approach to handling such load variations is auto-scaling, which adaptively adjusts computational resources (e.g., increasing or decreasing the number of GPUs in a cluster) auto_scaling .
-However, auto-scaling is primarily effective in cloud environments with abundant resources and typically operates on minute-to-hour timescales auto_scaling_timescale .
+A classic approach to handling such load variations is auto-scaling, which adaptively adjusts computational resources (e.g., increasing or decreasing the number of GPUs in a cluster) Catillo et al.
+(2023).
+However, auto-scaling is primarily effective in cloud environments with abundant resources and typically operates on minute-to-hour timescales Dogani et al.
+(2023).
 Consequently, it is inadequate for LLM serving systems, which must respond to sharp, second-level load spikes and scale throughput within millisecond-level intervals to meet SLO requirements.
@@
 Figure 1 presents the per-second p90 TPOT under three precision schemes: FP16, FP8, and the proposed dual-precision format, using the Llama 3.1 8B model on an H100 GPU with vLLM.
-The request patterns are extracted from a 60-second segment of Microsoft Azure’s LLM inference trace azure_llm_inference_trace , scaled down to 20% of the original load.
-The TPOT SLO threshold is set to 33 ms baseten_slo .
+The request patterns are extracted from a 60-second segment of Microsoft Azure’s LLM inference trace Azure (2024), scaled down to 20% of the original load.
+The TPOT SLO threshold is set to 33 ms Baseten (2025).
 As shown, the FP16 model exhibits more frequent TPOT spikes that exceed the TPOT SLO threshold compared to the FP8 model.
@@
 Handling Exception Layers.
-For layers containing weight values with magnitudes exceeding 1.75 (i.e., where the MSB of the exponent is non-zero), we retain them in FP16 and always perform FP16 computation.
+For layers containing weight values with magnitudes exceeding 1.75 (i.e., where the MSB of the exponent is possibly non-zero), we retain them in FP16 and always perform FP16 computation.
 This may reduce the benefits of dual-precision LLM serving, as the throughput advantage of FP8 mode becomes less pronounced.
@@
 FP16 reconstruction involves a large number of 8-bit bitwise operations.
-To reduce the associated overhead, we fuse four 8-bit operations into a single 32-bit instruction, following the approach in any_precision_llm ; fp6_llm .
+To reduce the associated overhead, we fuse four 8-bit operations into a single 32-bit instruction, following the approach in Park et al.
+(2024); Xia et al.
+(2024).
 Algorithm 1 illustrates how FP16 reconstruction is performed using merged bitwise operations.
@@
 For activation quantization, we apply per-token quantization for both NestedFP and the baseline.
-The evaluation covers six models: Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), Mistral Small (24B), Llama 3.1 70B, and DeepSeek-R1-Distill-Llama-70B deepseek_ref ; nemo_ref ; small_ref ; llama3_technical_report ; phi_ref .
-For benchmarks, we use three downstream tasks: Minerva Math, MMLU Pro, and BBH minerva ; bbh ; mmlupro .
-All models are integrated into the vLLM framework, and accuracy is measured using the LM Evaluation Harness lm-evaluation-harness .
+The evaluation covers six models: Llama 3.1 8B, Mistral Nemo (12B), Phi-4 (14B), Mistral Small (24B), Llama 3.1 70B, and DeepSeek-R1-Distill-Llama-70B AI (2024, 2025); DeepSeek-AI ; Meta (2024); Microsoft .
+For benchmarks, we use three downstream tasks: Minerva Math, MMLU Pro, and BBH Lewkowycz et al.
+(2022); Suzgun et al.
+(2023); Wang et al.
+(2024).
+All models are integrated into the vLLM framework, and accuracy is measured using the LM Evaluation Harness EleutherAI .
 As noted in Section 4.1, 8.75% of the linear layers in Phi-4 are not applicable to NestedFP and are therefore executed in FP16.
@@
 Across all models and benchmarks, the FP8 model in NestedFP achieves accuracy comparable to the baseline FP8 model, demonstrating the effectiveness of the new data format of NestedFP.
-Llama 3.1 8B Mistral Nemo Phi-4 Mistral Small Llama 3.1 70B DeepSeek-R1- Distill-Llama-70B FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N) Minerva Math 17.6 17.2 16.5 16.4 42.9 42.9 34.2 35.2 35.7 34.6 46.0 45.5 MMLU Pro 32.8 33.2 35.3 34.7 52.7 53.3 53.6 53.7 48.1 47.4 49.2 49.4 BBH 38.6 39.0 40.6 40.0 27.1 28.7 50.8 51.1 49.1 49.7 53.2 53.4
-Table 2: Accuracy of FP8 models in NestedFP across various downstream tasks, compared to the baseline FP8 model. FP8(B) denotes the baseline model, and FP8(N) denotes NestedFP.
+Minerva Math MMLU Pro BBH
+Model Size FP8(B) FP8(N) FP8(B) FP8(N) FP8(B) FP8(N)
+Llama 3.1 8B 17.6 17.2 32.8 33.2 38.6 39.0
+Mistral Nemo 12B 16.5 16.4 35.3 34.7 40.6 40.0
+Phi-4 14B 42.9 42.9 52.7 53.3 27.1 28.7
+Mistral Small 24B 34.2 35.2 53.6 53.7 50.8 51.1
+Llama 3.1 70B 35.7 34.6 48.1 47.4 49.1 49.7
+DeepSeek-R1-Distill-Llama 70B 46.0 45.5 49.2 49.4 53.2 53.4
+Table 2: Accuracy of FP8 models in NestedFP across various downstream tasks. FP8(B) denotes the baseline model, and FP8(N) denotes NestedFP.
 5.2 FP16 GEMM Kernel Performance
@@
 The results are compared against vanilla vLLM FP16 execution, which relies on PyTorch for invoking GEMM kernels.
-Again, the FP8-mode throughput of NestedFP is identical to baselines such as vanilla vLLM and is therefore omitted from evaluation.
+The FP8-mode throughput of NestedFP is excluded from the evaluation, as it does not require a separate kernel.
 Results.
 Figure 7 presents the results.
-NestedFP achieves FP16-mode throughput highly comparable to the baseline, with only minor overheads—5.04% (LLaMA 3.1–8B), 5.56% (Mistral Nemo), 4.76% (Phi-4), and 4.56% (Mistral Small) on average across batch sizes.
+NestedFP achieves FP16-mode throughput highly comparable to the baseline, with only minor overheads—5.04% (Llama 3.1 8B), 5.56% (Mistral Nemo), 4.76% (Phi-4), and 4.56% (Mistral Small) on average across batch sizes.
 The end-to-end throughput degradation is even lower than the kernel-level overheads reported in Section 5.2, since non-GEMM components also contribute to total execution time.
@@
 We measure p90 TTFT, p90 TPOT, and SLO attainment under two criteria: loose and tight.
-The loose SLO is defined as $\leq 10\times$ the average single-request latency, and the tight SLO is defined as $\leq 5\times$ the average single-request latency, following the methodology of a previous study sola_llm .
+The loose SLO is defined as $\leq 10\times$ the average single-request latency, and the tight SLO is defined as $\leq 5\times$ the average single-request latency, following the methodology of a previous study Hong et al.
+(2025).
 Results.
@@
 Appendix A Evaluation Details
-We conducted our evaluation using the Mistral Nemo Base 2407 and Mistral Small 24B Base 2501 models.
-For accuracy assessment, we employed the Math Verify Metric on the Minerva Math task, the Open LLM Leaderboard configuration (5-shot, multiple choice) for the MMLU Pro task, and a zero-shot evaluation setting for the BBH task.
-Kernel performance was measured using NVIDIA Nsight Compute.
-In order to get optimal kernel, we padded activation in the multiple of tile dimension $T_{m}$.
-During performance measurement, we flushed all caches using the cache-control option, but did not fix the GPU clock frequency using the clock-control option.
-Appendix B Extended Kernel Performance for NestedFP
-Figure 8: Performance comparison between our CUTLASS baseline and NestedFP kernel across 14 unique $(N,K)$ GEMM shapes.
-Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
-Figure 8 compares our method against the CUTLASS FP16 baseline across all 14 GEMM shapes for the four evaluated models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small.
-The matrix dimension $M$ is varied in increments of 32, ranging from 32 to 2048.
-Our results show that NestedFP consistently incurs only moderate overhead across all GEMM configurations, with an average performance difference of 6.1%.
-The per-graph average overhead ranges from 4.3% to 7.2%, demonstrating stable performance regardless of matrix shape.
-This consistency highlights the effectiveness of our kernel-level optimizations.
-Appendix C Extended End-to-End Throughput Evaluation for NestedFP
+Experiments were conducted on NVIDIA H100 GPU with vLLM 0.8.5 (V1 engine) and PyTorch 2.7.0+cu12.6.
+For accuracy assessment, we use the Math Verify metric on Minerva Math, the Open LLM Leaderboard configuration for MMLU Pro, and a zero-shot setting for BBH.
+The $M$ dimension is padded to multiples of $T_{m}$ only for kernel microbenchmarks and is disabled in the vLLM integration.
+Appendix B Extended Kernel Evaluation
+Figure 8: Comparison of kernel performance between the CUTLASS baseline and NestedFP.
+Figure 8 compares NestedFP with the CUTLASS FP16 baseline across 14 $(N,K)$ GEMM shapes for four models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small.
+We vary $M$ from 32 to 2048 in increments of 32.
+Overall, NestedFP incurs a moderate overhead, with an average of 6.38% relative to the FP16 baseline.
+Appendix C Extended End-to-End Throughput Evaluation
 (a) (32, 512)
@@
 (d) (1024, 32)
-Figure 9: Extended end-to-end throughput evaluation across different input/output token configurations.
-Each subplot shows throughput comparison for FP16 and NestedFP across four models (Llama 3.1 8B, Mistral Nemo, Phi-4, Mistral Small) under different request size settings: (a) 32 input, 512 output tokens (b) 1024 input, 512 output tokens (c) 32 input, 32 output tokens and (d) 1024 input, 32 output tokens.
-We extend our end-to-end evaluation to further demonstrate the robustness of NestedFP.
-Figure 9 presents results for four additional input/output configurations: (32, 512), (1024, 512), (32, 32), and (1024, 32).
-Across all settings, NestedFP exhibits minimal overhead compared to the FP16 baseline, with an average degradation of 5.52% and a maximum of 7.69%.
-On a per-model basis, the average overhead is 5.97% for LLaMA 3.1 (8B), 6.67% for Mistral-Nemo (12B), 5.51% for Phi-4 (14B), and 5.05% for Mistral-Small (24B).
-Our evaluation setup mirrors the configuration used in the main throughput experiments.
-Throughput was computed as the total number of generated tokens divided by the end-to-end latency for each test.
-All tests were conducted on a single NVIDIA H100 GPU using the vLLM 0.8.5 V1 engine with PyTorch 2.7.0+cu12.6 for both the baseline and our proposed NestedFP implementations.
-Chunked prefill, enabled by default in the vLLM V1 engine, was used throughout the evaluation.
-This feature segments the prefill phase into smaller chunks, which are processed concurrently with decoding tokens within the same batch.
-At each iteration, the GEMM operation shape $(M,N,K)$ is determined based on the number of tokens grouped in that step, with $M$ denoting the total token count.
-We set the maximum number of batched tokens to 8192 for the throughput test.
+Figure 9: FP16 inference throughput comparison across four models under different input/output token configurations.
+Figure 9 reports end-to-end throughput for (input, output) token pairs of (32, 512), (1024, 512), (32, 32), and (1024, 32).
+Across all settings, NestedFP achieves throughput comparable to the FP16 baseline.
+Chunked prefill was enabled, and the maximum number of batched tokens was set to 8192.
 Appendix D CUTLASS FP16 GEMM Kernel
-D.1 CUTLASS FP16 Baseline and PyTorch Kernel Comparison
-Figure 10: Performance comparison between PyTorch kernel and our CUTLASS baseline across 14 unique $(N,K)$ GEMM shapes.
-Each subplot represents a different shape configuration, with batch size $(M)$ on the $x$-axis and latency on the $y$-axis.
-To validate the robustness of our CUTLASS baseline, we compare its performance against the cuBLAS kernels employed by PyTorch 2.6.0 with CUDA 12.4.
-Using a weak baseline could underestimate the overhead introduced by NestedFP, leading to overly optimistic conclusions about its practical efficiency.
-Figure 10 provides a detailed comparison across 14 $(N,K)$ configurations, extracted from the linear layers of the four models evaluated in our main experiments—Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small.
-For each configuration, we sweep the batch size $M$ from 32 to 2048 in increments of 32, covering the typical range encountered in LLM inference workloads.
-Our results show that the FP16 baseline closely matches the performance of the cuBLAS implementation, with an average performance difference of just 1.8% across all 14 GEMM shapes.
-In 11 of the 14 configurations, our baseline achieves higher average performance than PyTorch, with performance differences reaching up to 3.3%.
-For the remaining three shapes—corresponding to smaller GEMMs—PyTorch slightly outperforms our implementation, with differences ranging from 0.3% to 0.9%, well within the bounds of normal GPU performance variability.
-These results confirm that our CUTLASS baseline is not only competitive but often exceeds production-level kernel performance, ensuring a reliable and fair comparison when evaluating NestedFP’s overhead.
+D.1 CUTLASS Baseline and PyTorch Kernel Comparison
+Figure 10: Comparison of kernel performance between CUTLASS baseline and PyTorch.
+Figure 10 compares kernel performance between the CUTLASS FP16 baseline and PyTorch 2.6.0 (CUDA 12.4).
+We evaluate 14 $(N,K)$ GEMM shapes from four models: Llama 3.1 8B, Mistral Nemo, Phi-4, and Mistral Small, sweeping $M$ from 32 to 2048 in increments of 32.
+Overall, the CUTLASS baseline achieves performance comparable to PyTorch.
 D.2 Kernel Search Space
-The kernel design space consists of two variants: non-cooperative and cooperative kernels.
-For non-cooperative kernels, we perform a grid search over tile dimensions with $T_{m}\in{16,32,64,128,256}$, $T_{n}\in{64,128,256}$, and $T_{k}\in{64,128,256}$.
-Cooperative kernels adopt larger tiles with $T_{n}\in{128,256}$, and configurations that fail to compile are excluded.
+We build upon CUTLASS 3.6.0.
+In CUTLASS, four warps are grouped into a warp group.
+We evaluate two variants: a non-cooperative kernel (one warp group for TMA and one for MMA) and a cooperative kernel (one warp group for TMA and two for MMA).
+For non-cooperative kernels, we grid-search tile dimensions with $T_{m}\in\{16,32,64,128,256\}$, $T_{n}\in\{64,128,256\}$, and $T_{k}\in\{64,128,256\}$.
+For cooperative kernels, we vary only $T_{n}$, setting $T_{n}\in\{128,256\}$.
+We exclude configurations that fail to compile.
 Non-cooperative kernels use a cluster shape of $(1,1,1)$, whereas cooperative kernels use $(2,1,1)$.
-We employ the persistent scheduler for non-cooperative kernels, and both the persistent and Stream-K [25] schedulers for cooperative kernels.
-Appendix E Broad Applicability of NestedFP
-Model GEMM1 GEMM2 GEMM3 GEMM4 Total
-CodeLlama 7B 96/96 32/32 64/64 31/32 223/224 (99.6%)
-CodeLlama 13B 120/120 40/40 80/80 37/40 277/280 (98.9%)
-Gemma 3 4B 207/264 64/88 123/176 34/34 429/563 (76.2%)
-Gemma 3 12B 249/306 78/102 151/204 48/48 527/661 (79.7%)
-Gemma 3 27B 291/348 92/116 179/232 62/62 625/759 (82.3%)
-Llama 3.1 8B 96/96 32/32 64/64 32/32 224/224 (100.0%)
-Llama 3.1 70B 224/240 80/80 141/160 78/80 523/560 (93.4%)
-Mistral Nemo 12B 120/120 40/40 80/80 40/40 280/280 (100.0%)
-Mistral Small 24B 120/120 40/40 80/80 40/40 280/280 (100.0%)
-Phi-3.5 Mini 26/32 31/32 31/32 24/32 112/128 (87.5%)
-Phi-4 14B 40/40 38/40 40/40 28/40 146/160 (91.2%)
-Qwen 3 8B 108/108 35/36 72/72 34/36 249/252 (98.8%)
-Qwen 3 14B 120/120 40/40 80/80 38/40 278/280 (99.3%)
-Qwen 3 32B 192/192 63/64 127/128 56/64 438/448 (97.8%)
-Table 4: Layer-wise applicability of NestedFP across models. Format X/Y indicates X applicable layers out of Y total layers for each GEMM type.
-Our scheme is broadly applicable to a wide range of LLMs beyond those evaluated in the main experiments.
-In Table 4, we analyze ten additional models that exhibit the key characteristic: the majority of their weight values have absolute magnitudes less than or equal to 1.75.
-For each model, we examine four distinct GEMM shapes corresponding to different linear layer types: GEMM1 (QKV projections), GEMM2 (output projections), GEMM3 (MLP gate/up projections), and GEMM4 (MLP down projections).
-Each entry is shown as X/Y, indicating that X out of Y layers of that type satisfy the weight magnitude constraint (absolute value $\leq$ 1.75).
-Our analysis indicates that most models exhibit high applicability rates.
-The Gemma 3 family shows the lowest rates (76.2%–82.3%), primarily due to multi-modal projection layers containing weights with significantly higher magnitudes (up to 26.25).
-Among the various GEMM types, GEMM4 (MLP down projections) tends to include more layers exceeding the threshold in models such as Phi-4 14B (70.0% applicability) and Qwen 3 32B (87.5%).
-While the majority of models have maximum weight values below 3.0, notable outliers include Llama 3.1 70B (maximum value of 93.0) and the Gemma 3 series (maximum value of 26.25).
-These extreme values are typically confined to a small number of layers and have limited impact on the overall applicability of our method.
-The consistently high applicability rates across diverse model families affirm the robustness and generalizability of our approach’s core assumption regarding weight distributions.
+We use the persistent scheduler for non-cooperative kernels, and consider both the persistent and Stream-K [25] schedulers for cooperative kernels.
+Appendix E Extended Applicability Analysis
+Model # Layers # Applicable Ratio (%)
+CodeLlama 7B 224 223 99.6
+CodeLlama 13B 280 277 98.9
+Gemma 3 4B 563 429 76.2
+Gemma 3 12B 661 527 79.7
+Gemma 3 27B 759 625 82.3
+Llama 3.1 8B 224 224 100.0
+Llama 3.1 70B 560 523 93.4
+Mistral Nemo 12B 280 280 100.0
+Mistral Small 24B 280 280 100.0
+Phi-3.5 Mini 128 112 87.5
+Phi-4 14B 160 146 91.2
+Qwen 3 8B 252 249 98.8
+Qwen 3 14B 280 278 99.3
+Qwen 3 32B 448 438 97.8
+Table 4: Layer-wise applicability of NestedFP across models.
+Table 4 presents an applicability analysis across a broader range of models.
+We report the total number of layers, the number of applicable layers satisfying the weight-magnitude constraint ($|w|\leq 1.75$), and the corresponding ratio.
+Overall, most models exhibit high applicability across layers.